### v5

# Algebroid functions on complex manifolds

Abstract.
We build a basic theory for algebroid functions on complex manifolds including several closely related topics such as uniformization problems, Nevanlinna theory, propagation of algebraic dependence and unicity problems, which is a development of the classical theory for algebroid functions on the complex plane. To the best of our knowledge, it is the first result for algebroid functions on a nontrivial complex manifold after H. Poincaré proposed the notion of algebroid functions.
Key words and phrases:
Algebroid functions; uniformization problems; second main theorem; Picard’s theorem; unicity.2010 Mathematics Subject Classification:
32H30; 32H25Contents
- 1 Introduction
- 2 Algebroid functions on Hermitian manifolds
- 3 Uniformization of algebroid functions on Hermitian manifolds
- 4 Nevanlinna theory of meromorphic functions on Kähler manifolds
- 5 Nevanlinna theory of algebroid functions on Kähler manifolds
- 6 Unicity problems of algebroid functions on Kähler manifolds
1. Introduction
1.1. Background
The notion of algebroid functions (see [22, 41]) introduced by H. Poincaré is a generalization of algebraic functions and meromorphic functions. Let us review that a -valued algebroid function on a (nonempty) domain (the complex plane) is defined by an irreducible algebraic equation
where are holomorphic functions with no common zeros on An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G. Darboux deemed that it is an important class of functions. To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection satisfying that it therefore realizes the uniformization of In fact, it leads to the famous uniformization problem for Riemann surfaces, which had been solved by Poincaré-Klein-Koebe (see [28, 35]) who showed the uniformization theorem which asserts that every Riemann surface is conformally equivalent to some quotient space where is one of the three canonical domains: or (the unit disc), while is a proper discrete group of Möbius automorphisms of
Even so, the most concerned problem on algebroid functions is their value distribution. About 1930s after R. Nevanlinna [33] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. began to study the value distribution theory of algebroid functions. For example, G. Rémoundos [36] proved a Picard’s theorem which asserts a nonconstant -valued algebroid function on can omit distinct values at most in Afterwards, G. Valiron [59] showed without details a five-value type theorem in unicity theory which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, involving investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc.. More details refer to H. Cartan [9], He-Xiao [22], Niino-Ozawa [29], J. Noguchi [30], M. Ozawa [34], M. Ru [37, 38], Sun-Gao [41], J. Suzuki [43], L. Selberg [45, 46], N. Toda [54, 55], E. Ullrich [58] and their references therein.
The algebroid function theory in one complex variable has received a great deal of achievements, however, it didn’t mature until the last 30-40 years. So far, there is very little literature discussing the algebroid functions in several complex variables, except part of contributions of Hu-Yang [21] to the theory on due to some technical difficulties such as the complexity of branching points and multiple points as well as uniformization of an algebroid function. We even know nothing about the algebroid functions on a nontrivial complex manifold. Given two nonconstant algebroid functions on a non-compact Hermitian manifold (referred to (1) in Section 2.1 for definition). We are inspired by the following closely related fundamental problems:
How to realize the uniformization of ?
How many Picard’s exceptional values can have at most?
How many values do have to share at least to guarantee ?
To solve the first problem, we construct a multi-leaf complex manifold via which is a generic multi-sheeted analytic covering of Then, we lift to a meromorphic function on via the natural projection with The strategy of solving the second problem is to establish the Nevanlinna theory of algebroid functions on a complete Kähler manifold under certain curvature condition. In order to establish the Second Main Theorem of our trick used is to lift to via Thereout, one can receive the defect relation of which derives the Picard’s theorem. The central idea of dealing with the last problem is the applications of Nevanlinna theory. In doing so, we generalize for the first time the propagation theory of algebraic dependence introduced by L. Smiley [47] to the algebroid functions. This is a technical tool to study unicity problems. Based on it, we obtain the sharp lower bound of the number of the shared values of guaranteeing
1.2. Main results
Our study of algebroid function theory covers four closely related aspects: uniformization problems, Nevanlinna theory, propagation theory of algebraic dependence and unicity problems. So, we shall state the main results of this paper from the following four parts in turn.
A. Uniformization problems
Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of some technical treatment, see Definition 2.15. Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (1) in a neighborhood of where is a branching point with order of If is a non-singular branching point of then we describe the asymptotic behavior of as approaches i.e., we show that can be expanded into the Pusieux series (see Theorem 2.14)
with around in a local holomorphic coordinate where are meromorphic functions in Pusieux series plays an important role in construction of the -leaf complex manifold of and so is further uniformized by lifting to That is:
Theorem A1.
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
Theorem A1 can be applied directly to the extension of the fields of meromorphic functions. Let be a generic -sheeted analytic covering of We denote by the fields of meromorphic functions on respectively. As a consequence of Theorem A1, we deduce that
Corollary A2.
is a -degree algebraic extension of
B. Nevanlinna theory
Nevanlinna theory in several complex variables has been developed rapidly, due to the outstanding work of a number of excellent scholars. For example, referring to Bott-Chern [8], Carlson-Griffiths-King [10, 20], H. Fujimoto [17], K. Kodaira [25], Noguchi-Winkelmann [31, 32], M. Ru [38], B. Shiffman [40], F. Sakai [42], W. Stoll [48, 49, 52] and Y. T. Siu [53], etc., and also refer the reader to A. Atsuji [1, 2, 3] and X. J. Dong [14, 15], etc.. In this paper, our central work is to establish the Nevanlinna theory of algebroid functions on a non-compact complete Kähler manifold based on estimates of Li-Yau [27] on heat kernels and so we can use which to derive the Picard’s theorem.
In this part and the following Parts B, C, D, is a non-compact complete Kähler manifold. Let be a -valued algebroid function on See definition in Section 5.3, the Nevanlinna’s functions of are defined under two kinds of different curvature conditions. To unify the form, we write
(b1) has non-negative Ricci curvature and maximal volume growth
Refer to [44, 56, 57], we see that such manifolds exist under this curvature condition. Let be the Ricci form of then we obtain:
Theorem B1 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Corollary B2 (Defect relation).
Assume the same conditions as in Theorem B1. Then
We obtain the following Picard’s theorem:
Corollary B3 (Picard Theorem).
Let be a transcendental -valued algebroid function on Then, can omit distinct values at most in
(b2) has non-positive sectional curvature
Let be the curvature function defined by (24) in Section 4.3. Under this curvature condition, we obtain:
Theorem B4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Corollary B5 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition
then
We obtain the following Picard’s theorem:
Corollary B6 (Picard Theorem).
Let be a nonconstant -valued algebroid function on satisfying the growth condition
Then, can omit distinct values at most in
C. Propagation theory of algebraic dependence
Propagation theory of algebraic dependence was introduced by L. Smiley [47]. Since it has important applications in unicity problems, then it aroused widespread concern. For example, the reader may refer to Y. Aihara [5, 6, 7], Dulock-Ru [12], S. Drouilhet [13], H. Fujimoto [18, 19], S. Ji [23, 24] and W. Stoll [51], etc.. The paper extends this theory to the algebroid functions via the Nevanlinna theory established in Part B.
Let be an effective divisor on where are irreducible hypersurfaces of Let be a positive integer (allowed to be ). Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ). Denote by the set of all nonconstant algebroid functions on Set
Let
be the set of all such that
In addition, if has non-positive sectional curvature, then we ask for
Let be the hyperplane line over We denote by the set of all indecomposable hypersurfaces in such that for some where and is the natural projection on the -th factor. Put
with and
Theorem C1.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Let be a positive integer. Set
Fix any Assume that for at least one Let be the set of all such that on Put
Corollary C2.
If then contains exactly one element.
D. Unicity problems
Five-value theorem is a well-known theorem in unicity theory, which states that two nonconstant meromorphic functions on are identical if they share five distinct values (ignoring multiplicities) in This theorem was later extended to algebroid functions on by G. Valiron [59] and further improved by Sun-Gao [41]. In our study, a sharp five-value type theorem for algebroid functions on a non-compact complete Kähler manifold is proved by applying the propagation theory of algebraic dependence established in Part C.
(d1) has non-negative Ricci curvature and maximal volume growth
Theorem D1.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
with then
Corollary D2 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
(d2) has non-positive sectional curvature
Theorem D3.
Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition
Let be distinct values in Assume that for all If satisfies
with then
Corollary D4 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition
If share distinct values ignoring multiplicities in then
2. Algebroid functions on Hermitian manifolds
2.1. Notion of algebroid functions
First of all, we shall illustrate how the notion of algebroid functions on can be generalized to non-compact Hermitian manifolds. This is the starting point of a theory on algebroid functions we would like to build in this paper.
Let be a -dimensional non-compact Hermitian manifold. We consider an irreducible algebraic equation of the unknown function over :
where are holomorphic functions locally defined on such that they define a meromorphic mapping
Let be the indeterminacy locus of which is an analytic set in of complex codimension not less than if or an empty set if So, is a holomorphic mapping of into
Remark 2.1.
The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for every satisfying that
where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on for
2.1.1. Meromorphic function elements
Set
Then, it gives that
Definition 2.2.
Let A pair is called a regular function element, if is holomorphic in a neighborhood of A pair is called a meromorphic function element, if is meromorphic in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
Theorem 2.3.
Let be a pair such that
Then there exists a unique regular function element subordinate to equation satisfying
Proof.
The theorem can be showed by employing implicit function theorem. Here, we present an alternative proof of complex analysis. Taking a such that (see Remark 2.1). We write in the form
where
The conditions mean that
By the continuation, there exist such that
for all satisfying and In addition, those are in In further, restrict to then we obtain
Employing Rouché’s theorem, for any fixed with and have the same number of zeros in the disc In fact, the number of zeros is 1, since in has only one zero in the disc when This implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that
Evidently, we see that is continuous on
Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking a local holomorphic coordinate around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
Note that
as That is,
Thus, is holomorphic in the holomorphic direction at This completes the proof. ∎
Definition 2.4.
A point is called a critical point, if either or has a multiple root as a polynomial in We use to denote the set of all critical points, called the critical set. A regular point means a point in where is called the regular set. Moreover, denote by the set of all points in such that has a multiple root as a polynomial in where is called the multiple set and a point in is called a multiple point.
For any the fundamental theorem of algebra asserts that
has exactly distinct complex roots. Therefore, Theorem 2.3 implies that
Corollary 2.5.
For any there exist exactly distinct regular function elements subordinate to equation
To see the common roots of and we define the resultant of and by a determinant:
where for According to Sylvester’s elimination method, we can see that if and only if or has a multiple root as a polynomial in i.e., This gives that is an analytic set of complex dimension Define the discriminant of by
Assume that splits into linear factors, say
A direct computation leads to
If then is a multiple root of as a polynomial in If we set and rewrite (1) as where
Since then one can see that is a zero of i.e., a pole of In fact, we will deduce from Theorem 2.23 (in Section 2.3) that if and only if To conclude, we see that Corollary 2.5 can be extended to meromorphic function elements as follows:
Corollary 2.6.
For any there exist exactly distinct meromorphic function elements subordinate to equation
2.1.2. Analytic continuation of meromorphic function elements
Recall that a meromorphic function element on is a pair where is a meromorphic function defined in a neighborhood of in As a matter of convenience, we write
for short without confusion. Instead of we sometimes take a geodesic ball neighborhood centered at with radius Let be any two meromorphic function elements on Define an equivalent relation as follows: if and only if
Set
We equip with a topology. A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of Apply analytic continuation, we can regard that
A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in Namely, In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying that In other words, it is defined to be A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It is not hard to check that it defines a topology of
Theorem 2.7.
is a Hausdorff space.
Proof.
Pick If then we can consider -neighborhoods of and of with Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation of meromorphic function elements.
Definition 2.8.
Let be two meromorphic function elements on Let be a curve with and We say that is an analytic continuation of along if
for any there is a meromorphic function element on
for any and any there exists such that for all and that is a direct analytic continuation of
Remark 2.9.
The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Indeed, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve. Thus, the analytic continuation is symmetric.
Theorem 2.10 (Analytic continuation).
Let be a curve. Then, a function element can continue analytically to along and attain to a unique function element
Proof.
Set
Then, we have We show Otherwise, we assume that By Corollary 2.6, there are exactly function elements at say that
Take small enough such that where Then there is a unique integer such that Notice that the domain of definition of is so for any one can redefine function elements Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Thus, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic (meromorphic) functions. ∎
Corollary 2.11.
Let be distinct function elements at respectively. Then can continue analytically to each other along any curve in
Theorem 2.12 (Connectivity).
Any two function elements in can continue analytically to each other along a curve in
Proof.
Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other. Let be the largest integer such that there are function elements in that can continue analytically to each other along a curve. Clearly, we have Without loss of generality, one may assume that these function elements are So, we can write
where are holomorphic functions in a neighborhood of By Vieta’s theorem, we obtain
It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds. If then we are able to determine another set of using the rest function elements However, this implies that equation (1) is reducible, a contradiction. ∎
Corollary 2.13.
is a path-connected space.
2.1.3. Properties of meromorphic function elements near multiple points
We discuss the properties of function elements near multiple points. Let be a pair satisfying that and
Take a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively. Make use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively. By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively. In the following, we will investigate the analytic continuation of a function element along a closed Jordan curve. Let and pick as a closed Jordan curve passing by around Without loss of generality, we extend analytically along started at When returns to we attain to a function If then one continues to extend analytically along Repeat this action, then we will see that cycles on and on periodically. Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (or multiplicity) of if Repeatedly, then is divided into finite groups. Let the number of groups. If the -th group has members, then we have
Without loss of generality, we consider the first group: Take a proper there is a biholomorphic mapping such that
where stands for the polydisc centered at with polyradius in and is a local holomorphic coordinate of in So, we have
Set and write Next, we shall show that is a single-valued analytic function on where is the unit disc centered at in the -plane. Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation. There exists a number such that on for Combining (1) with on we obtain
on If we obtain Thus, on and this bound does not depend on Whence, is bounded on This means that is holomorphic on and of the Taylor expansion
near Let be the set of all branching points of in Note that the value of is independent of if is sufficiently small. If is a non-singular branching point of i.e., does not lie in a crossing of components of then is smooth whenever is small enough. By and Riemann’s analytic continuation (because the boundedness of on and is properly chosen if necessary), there is a biholomorphic mapping with
such that can be expanded into the Pusieux series
near where are holomorphic functions in satisfying that for
According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a curve around in Observe the local expansion of is a -multiple -valued branching point (with order ) of
Assume that is not a singular branching point of Set we write
Similarly, we have
If then
Choose such that Put then
Thus, i.e., the case turns to the case Note that is a pole of order of
To conclude, for each we have confirmed that there are distinct single-valued meromorphic function elements subordinate to equation (1) if is not a branching point of If is a branching point of then we can have several multi-valued function elements subordinate to equation (1), say with where gives a branching point of with order In particular, if is a non-singular branching point of then there is the Pusieux expansion
near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also. By analytic continuation, equation (1) uniquely determines a -valued meromorphic function on outside a set of branching points of What is more, is separated into single-valued meromorphic components on a cut-manifold, say where is any real hypersurface of passing by all the branching points of such that is connected. Let denote the set of all the branching points of which is an analytic set contained in called the branching set of We obtain:
Theorem 2.14.
Let be a -valued function element subordinate to equation If is not a singular branching point of then can be expanded into the convergent Pusieux series
in a local holomorphic coordinate near such that and for where is the branching set in a local neighborhood of and are meromorphic functions in
Now, we introduce the definition of algebroid functions on a non-compact Hermitian manifold as follows.
Definition 2.15.
A -valued algebroid function on is defined by an irreducible algebraic equation
where are holomorphic functions locally defined on such that they define a meromorphic mapping from into
Remark 2.16.
By Definition 2.15, we see that the condition for is also necessary due to Vieta’s theorem. In other words, a -valued algebroid function on is defined equivalently by an irreducible algebraic equation
where are meromorphic functions on
2.2. Multi-leaf complex manifolds of algebroid functions
In this subsection, we will show how to construct a class of complex manifolds using algebroid functions. Such a class of manifolds have a remarkable feature of multifoliation. What’s important is that these multi-leaf complex manifolds are a way to the uniformization of algebroid functions. First of all, we deal with a simple situation, i.e., we equip with a complex structure such that is a non-compact complex manifold.
As a matter of convenience, a multi-valued function element is also called the multi-leaf function element; a (non-holomorphic) meromorphic function element is also called the polar function element; the polar function elements and multi-leaf function elements are called the singular function elements.
Let be an arbitrary function element. Let us consider a neighborhood of which is defined by such a set of all such that is a direct analytic continuation of satisfying i.e., we have Choose a proper if necessary, we may have a biholomorphic mapping such that in which is the unit ball centered at 0 in Define a natural mapping
via
Evidently, is a homeomorphism. Its inverse is
Set We show that is a complex atlas of Let It suffices to show the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Note that for each we have in a small neighborhood of i.e., A direct computation gives that
for Similarly, we obtain
for Hence, is a biholomorphic mapping. On the other hand, for any there corresponds to distinct function elements Therefore, we conclude that
Theorem 2.17.
is a non-compact -leaf complex manifold of complex dimension
Set
Corollary 2.18.
is a non-compact -leaf complex manifold of complex dimension which is a submanifold of
In what follows, we shall construct a -leaf complex manifold determined by all functions elements subordinate to equation (1). Let be any two function elements subordinate to equation (1). We define an equivalent relation as follows: if and only if
for some two single-valued components of respectively, where is the branching set of for Now, put the multi-leaf function elements in Set
Remark 2.19.
If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements. If then it can correspond to multi-leaf function elements in with where can be separated into single-valued components.
Definition 2.20.
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one can regard that
Let A neighborhood of is a set which contains the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
where is a neighborhood of In particular, a -neighborhood of (satisfying ), is defined by the set which contains the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
A subset is called an open set, if either is an empty set or every point is an inner point, namely, there exists a -neighborhood of such that If is a -leaf function element, then has leaves. In this case, we call a -leaf neighborhood of This gives a topology of and this topology contains the topology of
Theorem 2.21.
is a connected Hausdorff space.
Proof.
Hausdorff property of can be showed using the similar arguments as in the proof of Theorem 2.7. Next, we prove that is connected. If not, one can write where are two disjoint open subsets of Set It is evident that Notice that are two disjoint open subsets of since is an analytic set of complex codimension This means that is not connected, but it is a contradiction. ∎
Next, let us continue to endow with a complex structure such that is a complex manifold. Take any -leaf function element Assume that is a non-singular branching point of Apply Theorem 2.14 ( is properly chosen if necessary), there exists a biholomorphic mapping such that can be expanded into the Pusieux series
satisfying that and for
Let be a -leaf neighborhood of containing the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
Let be single-valued connected components of Define a mapping
by
where It is easy to check that is a homeomorphism and its inverse is given by
where is a positive integer such that Let be the set of all function elements such that is a singular branching point of some multi-valued component of and let be the set of all corresponding singular branching points. It is evident that is an analytic set of complex dimension not greater than Set
We will show that is a complex atlas of Let It is sufficient to prove the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Via a similar discussion, one only needs to deal with the situation where is a -leaf singular function element and is just a regular function element. For each there exists a single-valued component of such that in some neighborhood of i.e., It is clear that since all the function elements in are meromorphic. A direct computation leads to
where is a positive integer such that Hence, is a holomorphic mapping. On the other hand, its inverse is
where is the -th single-valued component of Evidently, is also a holomorphic mapping. Hence, is a non-compact -leaf complex manifold of complex dimension
Next, we consider the singular part Let be the natural projection by mapping a function element into Then, one sees that is a generic -sheeted analytic covering and
Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via Therefore, we obtain:
Theorem 2.22.
is a non-compact -leaf complex manifold of complex dimension called a complex manifold of a -valued algebroid function on
2.3. Zero and pole divisors of algebroid functions
Let be a -valued algebroid function on defined by equation As we have known in Section 2.1, is a pole of if and only if is a zero of However, it provides nothing about the relationship between the multiplicities of a pole of and a zero of The following theorem shows that they agree. Let be single-valued meromorphic components of As a matter of convenience, we write
Theorem 2.23.
We have
Proof.
Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors, and so we may assume that It yields that
where denotes the number of -valued points of a function at Without loss of generality, suppose that are all single-valued components such that which take at and suppose that are all single-valued components such that which take at Then, we have By Vieta’s theorem, it yields that
Combined this with (4), we see that reaches the maximal order of a pole of Since then we obtain
It follows that ∎
Theorem 2.24.
We have
Proof.
Set Apply the similar arguments as in the proof of Theorem 2.23, we can prove the theorem. ∎
3. Uniformization of algebroid functions on Hermitian manifolds
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing multi-leaf complex manifolds. Our idea is motivated by the work of W. Stoll [50] on algebroid reduction.
3.1. Algebroid reduction
Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, namely, it is a proper and surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex dimension It is evident that
is a proper, surjective, local biholomorphic mapping of sheet number and hence a -sheeted analytic covering. By Remmert [39] (see Andreotti-Stoll [4] also), the exceptional set
is an analytic set satisfying that and Moreover, and for Let the mapping degree of at is defined in such manner: if is a relatively compact neighborhood of such that then
which is independent of the choice of Note that and if and only if By W. Stoll [50], for any we have
Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection mapping. Let be the -fold symmetric tensor product of i.e.,
where
for any is called the symmetric tensor product of and is the -order symmetric group. It is evident that is a subspace of -fold tensor product space with Again, put
where for any Naturally, is embedded into by the identity mapping Thus, we obtain
Definition 3.1.
Let be a nonempty open subset of Let be a continuous mapping. Then, the algebroid reduction of via is a mapping defined by
Theorem 3.2 (Stoll, [50]).
is a continuous mapping. If is holomorphic, then is also holomorphic and extends to a holomorphic mapping
Definition 3.3.
Let be a meromorphic mapping with indeterminacy set Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
Theorem 3.4 (Stoll, [50]).
Let be a meromorphic mapping with indeterminacy locus Let be the algebroid reduction of via Fix any Assume that there exist a nonempty neighborhood of such that and a reduced representation If is a representation of on then the algebroid reduction of via is a representation of If is reduced, then is also reduced.
According to Theorem 3.4, we obtain the following commutative diagram:
Theorem 3.5 (Stoll, [50]).
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via
3.2. Uniformization problems
In what follows, we consider the case that For a convenience, we write and in short. Take a basis of Then for any it has a unique representation
for some numbers It gives a natural isomorphism defined by
which yields that
If then (6) defines an algebraic equation
Let be the maximal integer in such that Then, equation (7) reduces to Let be the roots of the equation. Apply Vieta’s theorem, it is immediate that
So, is decomposable, i.e.,
By this with (5)
To conclude, we obtain
Lemma 3.6.
Let be a differentiably non-degenerate and irreducible meromorphic mapping. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function which separates the fibers of such that is the algebroid reduction of via
Theorem 3.7 (Uniformization theorem).
Let be a -valued algebroid function on Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
Proof.
Let be defined by equation (1). It gives a differentiably non-degenerate and irreducible meromorphic mapping
Set with a basis Since (8), then we can write
in the sense of isomorphisms. Further, the Vieta’s theorem leads to
in the sense of isomorphisms. On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of (generic) sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
with
Let be -valued points of Then
By
we have the theorem proved. ∎
Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of (generic) sheet number via which is a proper surjective holomorphic mapping.
Corollary 3.8.
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
Let be a generic -sheeted analytic covering of a non-compact Hermitian manifold with branching set Denote by (resp. ) the field of all meromorphic functions on (resp. ).
We give the algebraic relation between and as follows.
Corollary 3.9.
is a -degree algebraic extension of
Proof.
For a meromorphic function on we see that is at most a -valued algebroid function with branching set on So, there exist meromorphic functions on such that
It implies that is a finite algebraic extension of of degree ∎
4. Nevanlinna theory of meromorphic functions on Kähler manifolds
Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to Kähler metric Recall that the Kähler form of is defined by
in a local holomorphic coordinate Moreover, we define the Ricci form of by
Invoking Wirtinger’s formula, we see that the Riemannian volume element of is
4.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point and denote by the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
where is the Dirac function on The harmonic measure on with respect to is defined by
where is the Riemannian volume element of and is the inward normal derivative on
Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of and here
so that
The characteristic function of for is defined by
Let The proximity function of for with respect to is defined by
where denotes the spherical distance on defined by
with
The counting function of for with respect to is defined by
We need the following Dynkin formula:
Lemma 4.1 (Dynkin formula, [3, 16]).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
We are to show the First Main Theorem:
Theorem 4.2 (First Main Theorem).
Assume that Then
Proof.
Apply Dynkin formula (see Lemma 4.1) to we obtain
Note that the first term on the left hand side equals and the right hand side equals
Hence, we have the theorem proved. ∎
4.2. Second Main Theorem for non-negative Ricci curvature
In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth, i.e.,
where is the volume of geodesic ball centered at a fixed reference point with radius Denote by the Riemannian distance function of from and by the area of which is the boundary of A global Green function of for with a pole at means that
where is the Dirac function on Put
By Li-Tam-Wang [26], there is a number independent of such that
as
If then there exists (see, e.g., [60, 61]) the unique minimal positive global Green function of for with a pole at According to the estimate due to Li-Yau [27] and (10), there exists a constant depending only on such that
Colding-Minicozzi II [11] (see Li-Tam-Wang [26] also) proved the following asymptotic behavior of :
Lemma 4.3.
For we have
Define the relatively compact -domain for by
Then, exhausts whenever is an increasing sequence with as Moreover, the boundary of is a submanifold of for almost every due to Sard’s theorem. Set
It is evident that gives the positive Green function of for with a pole at satisfying Dirichlet boundary condition. Also, denote by the harmonic measure on with respect to
If then curvature condition implies that is a parabolic Riemann surface, so each global Green function for must change sign. Again, since has maximal volume growth, we see that is conformally equivalent to i.e., there exists a biholomorphic mapping such that We equip with the pull-back metric induced from the standard Euclidean metric on Since gives a global Green function of for with a pole at then it corresponds to a global Green function of for with a pole at :
In this case, we define the relatively compact -domain for by
It is clear that
gives the Green function of for with a pole at satisfying Dirichlet boundary condition. A direct computation gives that the harmonic measure on with respect to is
where is the Riemannian volume element of
Let be a meromorphic function. Take in Section 4.1, then we have the Nevanlinna’s functions of :
4.2.1. Estimates of harmonic measures
We shall use the asymptotic behavior of (see Lemma 4.3) to derive both lower and upper bounds of the harmonic measure on
Theorem 4.4 (Gradient estimates of Green functions).
We have
If then
If then there exists a constant such that
holds for all sufficiently large
Proof.
If then is just the geodesic disc centered at with radius It means that for Thus, it follows from (12) that
where is the inward normal derivative on Hence, holds.
Next, we show that holds. Note that is continuous outside and the cut locus of has measure zero, one may assume without loss of generality that Since and as by L’Hôspital’s rule, then it yields from Lemma 4.3 that
for where stands for the inward normal derivative on On the other hand, Lemma 4.3 leads to
Hence, we see that approaches increasingly a geodesic ball centered at with radius as This also implies that
Combining the above, we obtain
Therefore, there exists a constant such that
holds for all sufficiently large ∎
Corollary 4.5 (Estimates of harmonic measures).
We have
If then
If then
for all sufficiently large In the above, is the Riemannian volume element of
Proof.
The corollary follows immediately from the following relation
4.2.2. Calculus Lemma
In order to establish the calculus lemma which plays an important role in the proof of the second main theorem later, we need to estimate the bounds of the Green function for in terms of integral forms. Set
Lemma 4.6 (Estimates of Green functions).
We have
If then
If then for any there exists sufficiently large such that for we have
Proof.
Theorem 4.7 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a positive constant and a subset with finite Lebesgue measure such that
holds for all outside
Proof.
If then is the geodesic disc centered at with radius It yields from Theorem 4.6 that
Set
Then
which yields from Corollary 4.5 that
On the other hand, applying Borel lemma to twice, then for any there exists a subset with finite Lebesgue measure such that
holds for all outside Combining (15) with (16), we have
holds for all outside where is a constant.
Next, we show the case where Invoking Lemma 4.6, for any there exists sufficiently large such that for we have
Notice that as due to Lemma 4.3, then there exists sufficiently large such that for we have
Hence, we conclude that
where
Set
A direct computation leads to
Furthermore, we have
According to Corollary 4.5, there exists sufficiently large such that
holds for all Thus, it follows from (18) that
Using Borel lemma twice, for any , we have
holds for all outside a subset with finite Lebesgue measure. Applying (19) and (20) and noting that as we see that there exists sufficiently large such that for all we have
Therefore, by this with (4.2.2)
holds for all outside where
is a constant. This proves the theorem. ∎
4.2.3. Logarithmic Derivative Lemma
To establish the second main theorem, we still need to prove a logarithmic derivative lemma. Let be the gradient operator on associated with the Kähler metric Let be a meromorphic function on The norm of the gradient of is defined by
in a local holomorphic coordinate where is the inverse of Define
where
Indeed, define for
where
It is not difficult to check that
On take a singular metric
which is satisfied with
To establish the Logarithmic Derivative Lemma, we require the following two lemmas.
Lemma 4.8.
We have
Proof.
A direct computation leads to
which yields from Fubini’s theorem that
Lemma 4.9.
Assume that Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Make use of the concavity of , we have
Thus, it yields from Theorem 4.7 that
for all outside The lemma is proved. ∎
Define
Theorem 4.10 (Logarithmic Derivative Lemma).
Let be a nonconstant meromorphic function on Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Note that
By Lemma 4.9, for any there exists a subset with finite Lebesgue measure such that
holds for all outside ∎
4.2.4. Second Main Theorem
The characteristic function of is defined by
Theorem 4.11 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Define a non-negative function on via
where
is a singular volume form on By we obtain
in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1), it yields that
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and with On each we write
where is a positive smooth function on Given a partition of the unity subordinate to Put
Again, set Then on we have
Fix a we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
It deduces that
Put Then, we obtain
on Therefore,
on Since is bounded on then it yields that
By this with Dynkin formula (see Lemma 4.1) and Theorem 4.10, for any there is a subset with finite Lebesgue measure such that
holds for all outside By (4.2.4) and (23), then the theorem holds. ∎
Let us consider the defect relation. Define the defect and the simple defect of with respect to respectively by
According to Theorem 4.2, we obtain Since has maximal volume growth, then whenever is nonconstant, as
Corollary 4.12 (Defect relation).
Assume the same conditions as in Theorem Then
4.3. Second Main Theorem for non-positive sectional curvature
In the subsection, we assume that has non-positive sectional curvature. Fix a point Review that denotes the geodesic ball centered at with radius By Sard’s lemma, the boundary of is a submanifold of for almost every Set
where is the pointwise lower bound of Ricci curvature defined by
It remains to use to denote the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and use to denote the harmonic measure on with respect to Let be a meromorphic function. Similarly, the Nevanlinna’s functions are defined by
Refer to Dong [14] (see Atsuji [3] also), we have the Second Main Theorem as follows:
Theorem 4.13 (Second Main Theorem).
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Similarly, we can define the defect and the simple defect of with respect to a value in
Corollary 4.14 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition
then
5. Nevanlinna theory of algebroid functions on Kähler manifolds
Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to its Kähler metric We adopt the same notations such as etc. defined in Section 4.
5.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definition in Section 4.1).
In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus Suppose The characteristic function of for is defined by
where
are called the proximity function, counting function of for respectively. For define
where
In order to unify the form, we write
We show the First Main Theorem:
Theorem 5.1 (First Main Theorem).
Assume that Then
Proof.
With the aid of Vieta’s theorem, we obtain
Using Dykin formula (see Lemma 4.1), then it yields that
From the definition, we see that the left hand side equals
On the other hand, Theorem 2.23 and Theorem 2.24 imply that the right hand side equals
Thus, we obtain
Set then
where
In particular, and Similarly, we have
which means that
Since
and
then we obtain
Put together the above, we conclude that
5.2. Estimates of branching divisors
Let be a -valued algebroid function on We use the same notations such as given in Section 5.1. Our aim is to give an estimate of the branching divisor of Denote by the -valued, -valued components of at any respectively. Then, we have namely, is a -order branching point of for and is hence a -order branching point of The branching divisor of associated to is defined by a divisor on satisfying that
The counting function of branching divisor of for is defined by
where
Recall that the discriminant of is given by
which is a -order determinant, see Section 2.1.
Lemma 5.2.
We have
in the sense of currents.
Proof.
Take any Let be the -valued, -valued components of at respectively, where Since the set of singular branching points of is an analytic set of complex dimension not greater than then one just needs to consider that is a non-singular branching point of By Theorem 2.14, we can write for
in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of One can note from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.23, we have
This implies the lemma holds. ∎
Set
Theorem 5.3 (Estimates of branching divisors).
We have
5.3. Second Main Theorem
Let be the -leaf complex manifold of -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.22. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on for that could be not positive-definite at points in the branching set of Using Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we show that Dykin formula (see Lemma 4.1) still holds under the pullback metric Set Pick a sequence of neighborhoods of such that
Equip with a complete Kähler metric Let Clearly, is a complete Kähler metric on and as Let be the Laplace-Beltrami operators of respectively. Then, weakly as Take a bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and by the harmonic measures of on with respect to respectively. It is evident that as and (weakly) uniformally on as due to
for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively. To unify the form, one uses to denote the volume elements of for respectively. Using Dynkin formula, it gives that
for a -class function with on outside a polar set of singularities at most. Hence, it yields immediately by letting that
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.11 and Theorem 4.13 still hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
5.3.1. with non-negative Ricci curvature
In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth. With the pullback metric has non-negative Ricci curvature and maximal volume growth. Let be the Kähler form of Take where is a bounded domain defined in Section 4.2. Let us recall the Green function for and the harmonic measure on defined in Section 4.2. Then, the characteristic function of is defined by
where
Indeed, define for
where
In order to unify the form, we write
The simple counting function of with respect to is defined by
Assume that We denote by the connected components of centered at respectively, with for Again, define
Then, it is trivial to see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and also is the harmonic measure on with respect to Set
where
Define for
and
If is not constant, then it follows from Theorem 4.11 that for distinct values there is a subset with finite Lebesgue measure such that
holds for and all outside
Theorem 5.4 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Proof.
Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (1), say where with and gives a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
On the other hand, we have and
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset with finite Lebesgue measure such that
holds for all outside Hence, we have the theorem proved. ∎
We continue to treat the defect relation. Define the defect and the simple defect of with respect to respectively by
According to Theorem 5.1, we obtain The condition of maximal volume growth gives that if is nonconstant, as
Corollary 5.5 (Defect relation).
Assume the same conditions as in Theorem Then
Corollary 5.6 (Picard Theorem).
Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in
5.3.2. with non-positive sectional curvature
In the subsection, we assume that has non-positive sectional curvature. Then, also has non-positive sectional curvature under the pullback metric Take where is the geodesic ball centered at with radius Recall the Green function for and the harmonic measure on in Section 4.3. Similarly, the characteristic function of is defined by
where
Indeed, for
where
In order to unify the form, we write
Also, define for
Assume that We denote by the connected components of centered at respectively, here for Again, define
We see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and is the harmonic measure on with respect to Put
with
where is the Ricci curvature tensor on Similarly, set
where
Also, define for
If is not constant, then it follows from Theorem 4.13 that for distinct values there is a subset with finite Lebesgue measure such that
holds for and all outside
Apply the similar arguments as in the proof of Theorem 5.4, we obtain:
Theorem 5.7 (Second Main Theorem).
Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
Set
Corollary 5.8 (Defect relation).
Assume the same conditions as in Theorem If satisfies the growth condition then
Corollary 5.9 (Picard Theorem).
Let be a nonconstant -valued algebroid function on satisfying the growth condition then can omit distinct values at most in
6. Unicity problems of algebroid functions on Kähler manifolds
In this section, is a non-compact complete Kähler manifold of complex dimension
6.1. Propagation of algebraic dependence
Let be the Fubini-Study form on Given a -valued algebroid function on and Define
Using Dynkin formula (see Lemma 4.1), it is immediate that
On the other hand, since
then we have
Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable. Given algebroid functions Put
Definition 6.1.
Let be an analytic subset of nonconstant algebroid functions on are said to be algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, we say that are -related on
Let be an effective divisor on and fix any positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
It is clear that Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ).
We introduce the notation as follows:
If has non-negative Ricci curvature and maximal volume growth, then we denote by
the set of all such that
If has non-positive sectional curvature, then we denote by
the set of all such that
and
Let stand for the hyperplane line bundle over i.e., is the holomorphic line bundle defined by a point in It gives a holomorphic line bundle over via
where is the natural projection on the -th factor with Let for For , define
with
for where (the first Chern class of ) is a fixed positive -form.
Lemma 6.2.
We have
Proof.
A direct computation leads to
Then, it follows that
Invoking Dynkin formula (see Lemma 4.1), it is therefore
This proves the lemma. ∎
Let be the set of all indecomposable hypersurfaces in such that for some Put
Lemma 6.3.
Assume that and for some Then
Proof.
Lemma 6.4.
Let be a -valued algebroid function in Then
Proof.
By we obtain
in the sense of currents, where and are understood as divisors. Thus, we obtain
By this with Theorem 5.4, we obtain
This proves the lemma. ∎
Set
with
We obtain a criteria for the propagation of algebraic dependence:
Theorem 6.5.
Let be a -valued algebrod function in for Assume that are -related on for some If then are -related on
Proof.
Put
Fix any -valued algebroid function We say that is generic with respect to if for at least one Assume that is generic with respect to Let be the set of all such that on Set
The following is a unicity theorem:
Theorem 6.6.
If then contains exactly one element.
Proof.
Let be a -valued algebroid function in It suffices to show that on Set
Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the hyperplane line bundle over If on then there is a holomorphic section of over such that and Taking then we obtain and On the other hand, implies that
where With the help of Theorem 6.5, we obtain which is a contradiction. This proves the theorem. ∎
6.2. Five-value type theorem
We shall apply the propagation theory of algebraic dependence established in Section 6.1 to the unicity problems of algebroid functions on non-compact complete Kähler manifolds under two kinds of different curvature conditions.
A1 has non-negative Ricci curvature and maximal volume growth
Under this curvature condition, we show that
Theorem 6.7.
Let be transcendental -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
In particular, the conclusions hold for nonconstant if
Proof.
Set and with Then
Making use of Theorem 6.6, we have holds. Similarly, if one sets for then we obtain
which also implies holds. ∎
Corollary 6.8 (Five-value type theorem).
Let be nonconstant -valued algebroid functions on If share distinct values ignoring multiplicities in then
B1 has non-positive sectional curvature
Let be defined by (24). Set
Apply the similar arguments as in the proof of Theorem 6.7, we obtain:
Theorem 6.9.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy the growth condition Then
Assume for all If then
Assume for all If then
Proof.
The theorem can be confirmed by carrying the same arguments as in the proof of Theorem 6.7. ∎
Corollary 6.10 (Five-value type theorem).
Let be nonconstant -valued algebroid functions on Assume that satisfy the growth condition If share distinct values ignoring multiplicities in then
Remark 6.11.
According to the well-known fact from the case for we see that the number of the shared values is sharp in Corollaries 6.8 and 6.10. However, Theorems 6.7 and 6.9 cannot yield the optimal number of the shared values if In doing so, we need to improve Theorems 6.7 and 6.9 so that one can derive the optimal number of the shared values.
A2 has non-negative Ricci curvature and maximal volume growth
To a general unicity problem for the value sharing ignoring multiplicities, we establish the following theorem:
Theorem 6.12.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
with then
Proof.
Assume on the contrary that The condition implies that
in the sense of currents, where and are understood as divisors. So, we are led to
where It yields from Lemma 6.3 that
Similarly,
Put together (6.2) and (6.2), it yields from Theorem 5.4 that
Therefore,
But, it contradicts with the assumption of The theorem is proved. ∎
Corollary 6.13.
Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Then
Assume for all If then
Assume for all If then
Proof.
Corollary 6.13 leads to a sharp five-value type theorem:
Corollary 6.14 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively. If share distinct values ignoring multiplicities in then
B2 has non-positive sectional curvature
Using the similar arguments as in the proof of Theorem 6.12, we obtain:
Theorem 6.15.
Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition Let be distinct values in Assume that for all If satisfies
with then
Corollary 6.16.
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition Let be distinct values in Then
Assume for all If then
Assume for all If then
Corollary 6.16 leads to a sharp five-value type theorem:
Corollary 6.17 (Five-value type theorem).
Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition If share distinct values ignoring multiplicities in then

### v9

@@
-Algebroid functions on complex manifolds
+Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds
 Abstract.
-We build a basic theory for algebroid functions on complex manifolds including several closely related topics such as uniformization problems, Nevanlinna theory, propagation of algebraic dependence and unicity problems, which is a development of the classical theory for algebroid functions on the complex plane.
-To the best of our knowledge, it is the first result for algebroid functions on a nontrivial complex manifold after H.
-Poincaré proposed the notion of algebroid functions.
+We generalize the classical Nevanlinna theory for algebroid functions on to a complete Kähler manifold with non-negative Ricci or non-positive sectional curvature.
+As an application, we give some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
 Key words and phrases:
-Algebroid functions; uniformization problems; second main theorem; Picard’s theorem; unicity.2010 Mathematics Subject Classification:
+algebroid functions; Nevanlinna theory; second main theorem; Picard type theorem.2010 Mathematics Subject Classification:
 32H30; 32H25Contents
@@
 - 2 Algebroid Functions on Hermitian Manifolds
-- 3 Uniformization of algebroid functions on Hermitian manifolds
-- 4 Nevanlinna theory of meromorphic functions on Kähler manifolds
-- 5 Nevanlinna theory of algebroid functions on Kähler manifolds
-- 6 Unicity problems of algebroid functions on Kähler manifolds
+- 3 Uniformization of Algebroid Functions
+- 4 Nevanlinna Theory of Meromorphic Functions
+- 5 Nevanlinna Theory of Algebroid Functions
+- 6 Unicity Problems of Algebroid Functions
 1. Introduction
-1.1. Background
+1.1. Motivation
 The notion of algebroid functions (see [16, 28]) introduced by H.
@@
 Darboux deemed that it is an important class of functions.
-To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection satisfying that it therefore realizes the uniformization of In fact, it leads to the famous uniformization problem for Riemann surfaces, which had been solved by Poincaré-Klein-Koebe (see [28, 35]) who showed the uniformization theorem which asserts that every Riemann surface is conformally equivalent to some quotient space where is one of the three canonical domains: or (the unit disc), while is a proper discrete group of Möbius automorphisms of
-Even so, the most concerned problem on algebroid functions is their value distribution.
-About 1930s after R.
+To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
+The most concerned problem on algebroid functions seems to be the value distribution.
+In the 1930s after R.
 Nevanlinna [22] founded two fundamental theorems concerning the value distribution theory of meromorphic functions in 1925, many famous scholars such as G.
@@
 For example, G.
-Rémoundos [36] proved a Picard’s theorem which asserts a nonconstant -valued algebroid function on can omit distinct values at most in Afterwards, G.
-Valiron [59] showed without details a five-value type theorem in unicity theory which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, involving investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc..
+Rémoundos [24] proved a Picard’s theorem which claims that a nonconstant -valued algebroid function on omits distinct values at most in Afterwards, G.
+Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, including investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc.
 More details refer to H.
@@
 Selberg [30, 31], N.
-Toda [54, 55], E.
-Ullrich [58] and their references therein.
-The algebroid function theory in one complex variable has received a great deal of achievements, however, it didn’t mature until the last 30-40 years.
-So far, there is very little literature discussing the algebroid functions in several complex variables, except part of contributions of Hu-Yang [21] to the theory on due to some technical difficulties such as the complexity of branching points and multiple points as well as uniformization of an algebroid function.
-We even know nothing about the algebroid functions on a nontrivial complex manifold.
-Given two nonconstant algebroid functions on a non-compact Hermitian manifold (referred to (1) in Section 2.1 for definition).
-We are inspired by the following closely related fundamental problems:
+Toda [37, 38] and E.
+Ullrich [39], etc.
+Algebroid function theory in one complex variable didn’t mature until the last 30-40 years.
+However, there is little literature on the algebroid functions in several complex variables, except for the work of Hu-Yang [15].
+Given two nonconstant algebroid functions on a non-compact Hermitian (Kähler) manifold .
+We are inspired by the following three fundamental problems:
 How to realize the uniformization of ?
 How many Picard’s exceptional values can have at most?
-How many values do have to share at least to guarantee ?
-To solve the first problem, we construct a multi-leaf complex manifold via which is a generic multi-sheeted analytic covering of Then, we lift to a meromorphic function on via the natural projection with The strategy of solving the second problem is to establish the Nevanlinna theory of algebroid functions on a complete Kähler manifold under certain curvature condition.
-In order to establish the Second Main Theorem of our trick used is to lift to via Thereout, one can receive the defect relation of which derives the Picard’s theorem.
-The central idea of dealing with the last problem is the applications of Nevanlinna theory.
-In doing so, we generalize for the first time the propagation theory of algebraic dependence introduced by L.
-Smiley [47] to the algebroid functions.
-This is a technical tool to study unicity problems.
-Based on it, we obtain the sharp lower bound of the number of the shared values of guaranteeing
+How many values should share at least to guarantee ?
+In this paper, we answer these three problems by generalizing the classical Nevanlinna theory for algebroid functions on to a non-compact complete Kähler manifold with non-negative Ricci or non-positive sectional curvature.
 1.2. Main Results
-Our study of algebroid function theory covers four closely related aspects: uniformization problems, Nevanlinna theory, propagation theory of algebraic dependence and unicity problems.
-So, we shall state the main results of this paper from the following four parts in turn.
-A. Uniformization problems
-Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of some technical treatment, see Definition 2.15.
-Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (1) in a neighborhood of where is a branching point with order of If is a non-singular branching point of then we describe the asymptotic behavior of as approaches i.e., we show that can be expanded into the Pusieux series (see Theorem 2.14)
-with around in a local holomorphic coordinate where are meromorphic functions in Pusieux series plays an important role in construction of the -leaf complex manifold of and so is further uniformized by lifting to That is:
-Theorem A1.
+A. Uniformization
+Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of a technical treatment (see Definition 2.15).
+Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (7) in a neighborhood of where is a branching point with order of If is a “non-singular” branching point of then we describe the asymptotic behavior of as approaches i.e., we prove that can be expanded into the Pusieux series (see Theorem 2.14)
+with around in local holomorphic coordinates where are meromorphic functions in With this Pusieux series, we show that
+Theorem A.
 Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
-Theorem A1 can be applied directly to the extension of the fields of meromorphic functions.
-Let be a generic -sheeted analytic covering of We denote by the fields of meromorphic functions on respectively.
-As a consequence of Theorem A1, we deduce that
-Corollary A2.
-is a -degree algebraic extension of
 B. Nevanlinna theory
-Nevanlinna theory in several complex variables has been developed rapidly, due to the outstanding work of a number of excellent scholars.
-For example, referring to Bott-Chern [8], Carlson-Griffiths-King [10, 20], H.
-Fujimoto [17], K.
-Kodaira [25], Noguchi-Winkelmann [31, 32], M.
-Ru [38], B.
-Shiffman [40], F.
-Sakai [42], W.
-Stoll [48, 49, 52] and Y.
-T.
-Siu [53], etc., and also refer the reader to A.
-Atsuji [1, 2, 3] and X.
-J.
-Dong [14, 15], etc..
-In this paper, our central work is to establish the Nevanlinna theory of algebroid functions on a non-compact complete Kähler manifold based on estimates of Li-Yau [27] on heat kernels and so we can use which to derive the Picard’s theorem.
-In this part and the following Parts B, C, D, is a non-compact complete Kähler manifold.
-Let be a -valued algebroid function on See definition in Section 5.3, the Nevanlinna’s functions of are defined under two kinds of different curvature conditions.
-To unify the form, we write
-(b1) has non-negative Ricci curvature and maximal volume growth
-Refer to [44, 56, 57], we see that such manifolds exist under this curvature condition. Let be the Ricci form of then we obtain:
+In this part and the following Parts C, D, we utilize to stand for a non-compact complete Kähler manifold.
+Let be a -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed reference point To unify the form, we write
+The simple defect of with respect to is defined by
+(b1) has non-negative Ricci curvature
+Assume that is non-parabolic. Set
+where denotes the Riemannian volume of the geodesic ball centered at with radius
 Theorem B1 (Second Main Theorem).
-Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
-holds for all outside
+Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is defined by
+When we have Hence, Theorem B1 covers the classical second main theorem for with (see Corollary 5.5).
 Corollary B2 (Defect Relation).
 Assume the same conditions as in Theorem B1. Then
-We obtain the following Picard’s theorem:
-Corollary B3 (Picard Theorem).
-Let be a transcendental -valued algebroid function on Then, can omit distinct values at most in
+if one of the following conditions is satisfied
+satisfies the volume growth condition
+where is defined by
+is of non-polynomial type growth, i.e., satisfies the growth condition
+We have as due to
+The condition (2) is relaxed. For example, it holds when with and holds when with or and
+Corollary B3 (Picard Type Theorem).
+Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in if one of the conditions in Corollary is satisfied.
 (b2) has non-positive sectional curvature
-Let be the curvature function defined by (24) in Section 4.3. Under this curvature condition, we obtain:
+Assume that is non-positively curved satisfying
+for constants with that either or where are the sectional curvature and Ricci curvature of respectively. Set
 Theorem B4 (Second Main Theorem).
-Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
-holds for all outside
+Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is defined by
+Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
 Corollary B5 (Defect Relation).
-Assume the same conditions as in Theorem If satisfies the growth condition
+Assume the same conditions as in Theorem B4. If
 then
-We obtain the following Picard’s theorem:
-Corollary B6 (Picard Theorem).
-Let be a nonconstant -valued algebroid function on satisfying the growth condition
-Then, can omit distinct values at most in
-C. Propagation theory of algebraic dependence
-Propagation theory of algebraic dependence was introduced by L.
-Smiley [47].
-Since it has important applications in unicity problems, then it aroused widespread concern.
-For example, the reader may refer to Y.
-Aihara [5, 6, 7], Dulock-Ru [12], S.
+Corollary B6 (Picard Type Theorem).
+Let be a nonconstant -valued algebroid function on satisfying Then, can omit distinct values at most in
+C. Propagation of algebraic dependence
+Propagation problem for algebraic dependence of meromorphic mappings was first studied by L.
+Smiley [32] and further studied by Y.
+Aihara [3, 4, 5], S.
 Drouilhet [9], H.
-Fujimoto [18, 19], S.
-Ji [23, 24] and W.
-Stoll [51], etc..
-The paper extends this theory to the algebroid functions via the Nevanlinna theory established in Part B.
+Fujimoto [13, 14], Dulock-Ru [10] and W.
+Stoll [36], etc.
+Assume that is either non-parabolic with non-negative Ricci curvature, or has non-positive sectional curvature satisfying (3).
 Let be an effective divisor on where are irreducible hypersurfaces of Let be a positive integer (allowed to be ). Set
@@
 be the set of all such that
-In addition, if has non-positive sectional curvature, then we ask for
+and (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has non-positive sectional curvature satisfying (3).
 Let be the hyperplane line over We denote by the set of all indecomposable hypersurfaces in such that for some where and is the natural projection on the -th factor. Put
@@
 D. Unicity problems
-Five-value theorem is a well-known theorem in unicity theory, which states that two nonconstant meromorphic functions on are identical if they share five distinct values (ignoring multiplicities) in This theorem was later extended to algebroid functions on by G.
-Valiron [59] and further improved by Sun-Gao [41].
-In our study, a sharp five-value type theorem for algebroid functions on a non-compact complete Kähler manifold is proved by applying the propagation theory of algebraic dependence established in Part C.
-(d1) has non-negative Ricci curvature and maximal volume growth
+(d1) has non-negative Ricci curvature
+Assume that is non-parabolic satisfying (2). Let be a positive integer. Consider the following condition
+with
 Theorem D1.
-Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
-with then
+Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
 Corollary D2 (Five-Value Type Theorem).
@@
 (d2) has non-positive sectional curvature
+Assume that is non-positively curved satisfying (3).
 Theorem D3.
-Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition
-Let be distinct values in Assume that for all If satisfies
-with then
+Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
 Corollary D4 (Five-Value Type Theorem).
-Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition
-If share distinct values ignoring multiplicities in then
+Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then
 2. Algebroid Functions on Hermitian Manifolds
 2.1. Notion of Algebroid Functions
-First of all, we shall illustrate how the notion of algebroid functions on can be generalized to non-compact Hermitian manifolds.
-This is the starting point of a theory on algebroid functions we would like to build in this paper.
-Let be a -dimensional non-compact Hermitian manifold. We consider an irreducible algebraic equation of the unknown function over :
-where are holomorphic functions locally defined on such that they define a meromorphic mapping
+First of all, we would like to extend the notion of algebroid functions to a non-compact Hermitian manifold.
+Let be a -dimensional non-compact Hermitian manifold.
+Consider an irreducible algebraic equation in on :
+where are holomorphic functions locally defined on such that they well define a meromorphic mapping
 Let be the indeterminacy locus of which is an analytic set in of complex codimension not less than if or an empty set if So, is a holomorphic mapping of into
 Remark 2.1.
-The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for every satisfying that
+The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for all satisfying that
 where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on for
@@
 Set
-Then, it gives that
+It gives that
 Definition 2.2.
@@
 Employing Rouché’s theorem, for any fixed with and have the same number of zeros in the disc In fact, the number of zeros is 1, since in has only one zero in the disc when This implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that
-Evidently, we see that is continuous on
-Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking a local holomorphic coordinate around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
+Evidently, is continuous on
+Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking local holomorphic coordinates around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
 Note that
@@
 For any there exist exactly distinct regular function elements subordinate to equation
-To see the common roots of and we define the resultant of and by a determinant:
+To see the common roots of and we define the resultant of and by a -order determinant:
 where for According to Sylvester’s elimination method, we can see that if and only if or has a multiple root as a polynomial in i.e., This gives that is an analytic set of complex dimension Define the discriminant of by
@@
 A direct computation leads to
-If then is a multiple root of as a polynomial in If we set and rewrite (1) as where
+If then has a multiple root as a polynomial in If we set and rewrite (7) as where
 Since then one can see that is a zero of i.e., a pole of In fact, we will deduce from Theorem 2.22 (in Section 2.3) that if and only if To conclude, we see that Corollary 2.5 can be extended to meromorphic function elements as follows:
@@
 We equip with a topology.
-A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of Apply analytic continuation, we can regard that
+A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we can regard that
 A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in Namely, In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying that In other words, it is defined to be A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It is not hard to check that it defines a topology of
@@
 Proof.
-Pick If then we can consider -neighborhoods of and of with Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
-Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation of meromorphic function elements.
+Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
+Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation) of meromorphic function elements.
 Definition 2.8.
@@
 Remark 2.9.
-The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Indeed, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve.
+The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve.
 Thus, the analytic continuation is symmetric.
@@
 Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other.
-Let be the largest integer such that there are function elements in that can continue analytically to each other along a curve.
-Clearly, we have Without loss of generality, one may assume that these function elements are So, we can write
+Let be the largest integer such that there are function elements in which can continue analytically to each other along a curve.
+Clearly, we have Without loss of generality, we shall assume that these function elements are So, we can write
 where are holomorphic functions in a neighborhood of By Vieta’s theorem, we obtain
 It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds.
-If then we are able to determine another set of using the rest function elements However, this implies that equation (1) is reducible, a contradiction.
+If then we are able to determine another set via the rest function elements However, it implies that equation (7) is reducible, which is a contradiction.
 ∎
@@
 2.1.3. Properties of meromorphic function elements near multiple points
-We discuss the properties of function elements near multiple points. Let be a pair satisfying that and
-Take a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively.
-Make use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively.
+We discuss the local properties of function elements near multiple points. Let be a pair satisfying that and
+Taking a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively.
+By the use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively.
 By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively.
-In the following, we will investigate the analytic continuation of a function element along a closed Jordan curve.
-Let and pick as a closed Jordan curve passing by around Without loss of generality, we extend analytically along started at When returns to we attain to a function If then one continues to extend analytically along Repeat this action, then we will see that cycles on and on periodically.
-Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (or multiplicity) of if Repeatedly, then is divided into finite groups.
-Let the number of groups.
+In the following, we discuss the analytic continuation of a function element along a closed Jordan curve surrounding a multiple point.
+Take Pick as a closed Jordan curve passing by around Without loss of generality, we only extend analytically along starting from When returns to from along one will attain to a function If then one continues to extend analytically along Repeating this action, we see finally that cycles on and on periodically.
+Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (multiplicity) of if Repeatedly, then it is not hard to verify that can be divided into several groups.
+Let be the number of groups.
 If the -th group has members, then we have
-Without loss of generality, we consider the first group: Take a proper there is a biholomorphic mapping such that
+Without loss of generality, we consider the first group: Take a proper we have a biholomorphic mapping such that
 where stands for the polydisc centered at with polyradius in and is a local holomorphic coordinate of in So, we have
@@
 Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation.
-There exists a number such that on for Combining (1) with on we obtain
+There exists a number such that on for Combining (7) with that on we obtain
 on If we obtain Thus, on and this bound does not depend on Whence, is bounded on This means that is holomorphic on and of the Taylor expansion
@@
 near where are holomorphic functions in satisfying that for
-According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a curve around in Observe the local expansion of is a -multiple -valued branching point (with order ) of
-Assume that is not a singular branching point of Set we write
+According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a closed curve around in By a local expansion of is a -multiple -valued branching point (with order ) of
+Assume that is not a singular branching point of Set and write
 Similarly, we have
@@
 Choose such that Put then
-Thus, i.e., the case turns to the case Note that is a pole of order of
-To conclude, for each we have confirmed that there are distinct single-valued meromorphic function elements subordinate to equation (1) if is not a branching point of If is a branching point of then we can have several multi-valued function elements subordinate to equation (1), say with where gives a branching point of with order In particular, if is a non-singular branching point of then there is the Pusieux expansion
+Thus, i.e., the case turns to the case Note that is a pole with order of
+To conclude, for each we have confirmed that there are distinct (single-valued) meromorphic function elements subordinate to equation (7), if is not a branching point.
+If is a branching point, then one shows that there are several multi-valued function elements (which contain single-valued function elements) subordinate to equation (7), say where is a -order branching point of satisfying At this time, we call a -order branching point of If is a non-singular branching point of then we have the convergent Pusieux expansion
 near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also.
-By analytic continuation, equation (1) uniquely determines a -valued meromorphic function on outside a set of branching points of What is more, is separated into single-valued meromorphic components on a cut-manifold, say where is any real hypersurface of passing by all the branching points of such that is connected.
-Let denote the set of all the branching points of which is an analytic set contained in called the branching set of We obtain:
+By the analytic continuation, equation (7) uniquely defines a -valued meromorphic function on outside a set of branching points of and is separated into single-valued meromorphic components on Let denote the set of all branching points of which is an analytic set contained in called the branching set of As a result, we conclude that
 Theorem 2.14.
 Let be a -valued function element subordinate to equation If is not a singular branching point of then can be expanded into the convergent Pusieux series
-in a local holomorphic coordinate near such that and for where is the branching set in a local neighborhood of and are meromorphic functions in
-Now, we introduce the definition of algebroid functions on a non-compact Hermitian manifold as follows.
+in a local holomorphic coordinate near such that =0 and for where is the branching set in a local neighborhood of and are meromorphic functions in
+We conclude the definition of algebroid functions on as follows.
 Definition 2.15.
 A -valued algebroid function on is defined by an irreducible algebraic equation
-where are holomorphic functions locally defined on such that they define a meromorphic mapping from into
-Remark 2.16.
-By Definition 2.15, we see that the condition for is also necessary due to Vieta’s theorem. In other words, a -valued algebroid function on is defined equivalently by an irreducible algebraic equation
-where are meromorphic functions on
+where are holomorphic functions locally defined on such that they well define a meromorphic mapping from into
+Let denote distinct (single-valued) meromorphic components of Given any -leaf function element of with If denote the single-valued components of then there exists a permutation of such that on In other words, defines a -leaf function element on Thus, any function element of can be represented locally by one or several single-valued components of
 2.2. Multi-leaf Complex Manifolds of Algebroid Functions
-In this subsection, we will show how to construct a class of complex manifolds using algebroid functions.
-Such a class of manifolds have a remarkable feature of multifoliation.
-What’s important is that these multi-leaf complex manifolds are a way to the uniformization of algebroid functions.
-First of all, we deal with a simple situation, i.e., we equip with a complex structure such that is a non-compact complex manifold.
-As a matter of convenience, a multi-valued function element is also called the multi-leaf function element; a (non-holomorphic) meromorphic function element is also called the polar function element; the polar function elements and multi-leaf function elements are called the singular function elements.
-Let be an arbitrary function element.
-Let us consider a neighborhood of which is defined by such a set of all such that is a direct analytic continuation of satisfying i.e., we have Choose a proper if necessary, we may have a biholomorphic mapping such that in which is the unit ball centered at 0 in Define a natural mapping
+Let be a function element.
+A neighborhood of is defined as a set of all such that is a direct analytic continuation of satisfying i.e., Taking a proper if necessary, there exists a biholomorphic mapping such that where is the unit ball centered at the coordinate origin in Define a natural mapping
 via
@@
 We define an equivalent relation as follows: if and only if
-for some two single-valued components of respectively, where is the branching set of for Now, put the multi-leaf function elements in Set
+for some two single-valued components of respectively, where is the branching set of for Set
 Remark 2.18.
 If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements.
-If then it can correspond to multi-leaf function elements in with where can be separated into single-valued components.
+If then it can correspond to multi-leaf function elements in where is a -order branching point of satisfying In other words, can be separated into single-valued components for
 Definition 2.19.
@@
 all satisfying
-Let be single-valued connected components of Define a mapping
-by
-where It is easy to check that is a homeomorphism and its inverse is given by
+Let be single-valued connected components of Define a mapping by
+where Note that is a homeomorphism and its inverse is
 where is a positive integer such that Let be the set of all function elements such that is a singular branching point of some multi-valued component of and let be the set of all corresponding singular branching points.
 It is evident that is an analytic set of complex dimension not greater than Set
-We will show that is a complex atlas of Let It is sufficient to prove the transition mapping
+We will show that is a complex atlas of Suppose It is sufficient to prove the transition mapping
 is a biholomorphic mapping.
 This is automatically a homeomorphism.
-Via a similar discussion, one only needs to deal with the situation where is a -leaf singular function element and is just a regular function element.
-For each there exists a single-valued component of such that in some neighborhood of i.e., It is clear that since all the function elements in are meromorphic.
+Via a similar discussion, one only needs to deal with the situation where is a -leaf meromorphic function element and is a regular function element.
+For every there exists a single-valued component of such that in some neighborhood of i.e., It is clear that because all function elements in are meromorphic.
 A direct computation leads to
@@
 where is the -th single-valued component of Evidently, is also a holomorphic mapping. Hence, is a non-compact -leaf complex manifold of complex dimension
-Next, we consider the singular part Let be the natural projection by mapping a function element into Then, one sees that is a generic -sheeted analytic covering and
-Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via Therefore, we obtain:
+Finally, we treat the singular part Let be the natural projection by mapping a function element into Then, we note that is a generic -sheeted analytic covering such that
+Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via To conclude, we obtain:
 Theorem 2.21.
-is a non-compact -leaf complex manifold of complex dimension called a complex manifold of a -valued algebroid function on
+is a non-compact -leaf complex manifold of complex dimension which is called the complex manifold of
 2.3. Zero and Pole Divisors of Algebroid Functions
-Let be a -valued algebroid function on defined by equation As we have known in Section 2.1, is a pole of if and only if is a zero of However, it provides nothing about the relationship between the multiplicities of a pole of and a zero of The following theorem shows that they agree.
-Let be single-valued meromorphic components of As a matter of convenience, we write
+Let be a -valued algebroid function on defined by equation Let be single-valued meromorphic components of As a matter of convenience, we write
 Theorem 2.22.
@@
 Proof.
-Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors, and so we may assume that It yields that
+Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors. Thus, we assume that It yields that
 where denotes the number of -valued points of a function at Without loss of generality, suppose that are all single-valued components such that which take at and suppose that are all single-valued components such that which take at Then, we have By Vieta’s theorem, it yields that
-Combined this with (4), we see that reaches the maximal order of a pole of Since then we obtain
-It follows that ∎
+Combining this with (10), we see that reaches the maximal order of a pole of Since then we obtain
+which follows that ∎
 Theorem 2.23.
@@
 Set Apply the similar arguments as in the proof of Theorem 2.22, we can prove the theorem. ∎
-3. Uniformization of algebroid functions on Hermitian manifolds
+3. Uniformization of Algebroid Functions
 This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold.
-As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing multi-leaf complex manifolds.
-Our idea is motivated by the work of W.
+As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold.
+Our main idea comes from the excellent work of W.
 Stoll [35] on algebroid reduction.
@@
 Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via
-3.2. Uniformization problems
+3.2. Uniformization
 In what follows, we consider the case that For a convenience, we write and in short. Take a basis of Then for any it has a unique representation
 for some numbers It gives a natural isomorphism defined by
-which yields that
-If then (6) defines an algebraic equation
+which yields that If then (12) defines an algebraic equation
 Let be the maximal integer in such that Then, equation (13) reduces to Let be the roots of the equation. Apply Vieta’s theorem, it is immediate that
 So, is decomposable, i.e.,
-By this with (5)
-To conclude, we obtain
+By this with (11), we have To conclude, we obtain
 Lemma 3.6.
@@
 Let be defined by equation (7). It gives a differentiably non-degenerate and irreducible meromorphic mapping
-Set with a basis Since (8), then we can write
+Set with a basis By (14), we can write
 in the sense of isomorphisms. Further, the Vieta’s theorem leads to
 in the sense of isomorphisms.
-On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of (generic) sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
+On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
 with
@@
 we have the theorem proved. ∎
-Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of (generic) sheet number via which is a proper surjective holomorphic mapping.
+Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of generic sheet number via which is a proper surjective holomorphic mapping.
 Corollary 3.8.
 Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
-Let be a generic -sheeted analytic covering of a non-compact Hermitian manifold with branching set Denote by (resp. ) the field of all meromorphic functions on (resp. ).
-We give the algebraic relation between and as follows.
-Corollary 3.9.
-is a -degree algebraic extension of
-Proof.
-For a meromorphic function on we see that is at most a -valued algebroid function with branching set on So, there exist meromorphic functions on such that
-It implies that is a finite algebraic extension of of degree ∎
-4. Nevanlinna theory of meromorphic functions on Kähler manifolds
-Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to Kähler metric Recall that the Kähler form of is defined by
-in a local holomorphic coordinate Moreover, we define the Ricci form of by
-Invoking Wirtinger’s formula, we see that the Riemannian volume element of is
+4. Nevanlinna Theory of Meromorphic Functions
+Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
+in local holomorphic coordinates Moreover, we define the Ricci form of by
 4.1. First Main Theorem
 Let be an arbitrary relatively compact domain with smooth boundary in Fix a point and denote by the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
-where is the Dirac function on The harmonic measure on with respect to is defined by
+where is the Dirac’s delta function on with a pole at The harmonic measure on with respect to is defined by
 where is the Riemannian volume element of and is the inward normal derivative on
-Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of and here
-so that
+Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
 The characteristic function of for is defined by
@@
 The counting function of for with respect to is defined by
-We need the following Dynkin formula:
-Lemma 4.1 (Dynkin formula, [3, 16]).
+Lemma 4.1 (Dynkin formula, [1, 11, 12]).
 Let be a -class function on outside a polar set of singularities at most. Assume that Then
-We are to show the First Main Theorem:
 Theorem 4.2 (First Main Theorem).
@@
 Note that the first term on the left hand side equals and the right hand side equals
-Hence, we have the theorem proved. ∎
 4.2. Second Main Theorem for Non-negative Ricci Curvature
-In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth, i.e.,
-where is the volume of geodesic ball centered at a fixed reference point with radius Denote by the Riemannian distance function of from and by the area of which is the boundary of A global Green function of for with a pole at means that
-where is the Dirac function on Put
-By Li-Tam-Wang [26], there is a number independent of such that
-as
-If then there exists (see, e.g., [60, 61]) the unique minimal positive global Green function of for with a pole at According to the estimate due to Li-Yau [27] and (10), there exists a constant depending only on such that
-Colding-Minicozzi II [11] (see Li-Tam-Wang [26] also) proved the following asymptotic behavior of :
+Assume that is non-parabolic with non-negative Ricci Curvature.
+Fix a reference point and denote by the geodesic ball centered at with radius in Denote by the Riemannian volume of The non-parabolicity of means that
+Thus, we have the unique minimal positive global Green function of for which can be written as
+where is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
+holds for all
+4.2.1. Construction of and Nevanlinna’s functions
+For define
+Since and then we conclude that is a precompact domain containing satisfying that and In fact, the family exhausts i.e., for any sequence with we have
+So, the boundary of can be formulated as
+By Sard’s theorem, is a submanifold of for almost all
+Set
+Note that defines the Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
+Let denote the harmonic measure on with respect to defined by
+where is the inward normal derivative on is the Riemannian area element of
+Let be a meromorphic function. Take in Section 4.1, we have the Nevanlinna’s functions of :
+4.2.2. Gradient estimates of Green functions
 Lemma 4.3.
 For we have
-Define the relatively compact -domain for by
-Then, exhausts whenever is an increasing sequence with as Moreover, the boundary of is a submanifold of for almost every due to Sard’s theorem. Set
-It is evident that gives the positive Green function of for with a pole at satisfying Dirichlet boundary condition. Also, denote by the harmonic measure on with respect to
-If then curvature condition implies that is a parabolic Riemann surface, so each global Green function for must change sign.
-Again, since has maximal volume growth, we see that is conformally equivalent to i.e., there exists a biholomorphic mapping such that We equip with the pull-back metric induced from the standard Euclidean metric on Since gives a global Green function of for with a pole at then it corresponds to a global Green function of for with a pole at :
-In this case, we define the relatively compact -domain for by
-It is clear that
-gives the Green function of for with a pole at satisfying Dirichlet boundary condition. A direct computation gives that the harmonic measure on with respect to is
-where is the Riemannian volume element of
-Let be a meromorphic function. Take in Section 4.1, then we have the Nevanlinna’s functions of :
-4.2.1. Estimates of harmonic measures
-We shall use the asymptotic behavior of (see Lemma 4.3) to derive both lower and upper bounds of the harmonic measure on
-Theorem 4.4 (Gradient estimates of Green functions).
-We have
-If then
-If then there exists a constant such that
-holds for all sufficiently large
+holds for all where is given by
 Proof.
-If then is just the geodesic disc centered at with radius It means that for Thus, it follows from (12) that
-where is the inward normal derivative on Hence, holds.
-Next, we show that holds.
-Note that is continuous outside and the cut locus of has measure zero, one may assume without loss of generality that Since and as by L’Hôspital’s rule, then it yields from Lemma 4.3 that
-for where stands for the inward normal derivative on On the other hand, Lemma 4.3 leads to
-Hence, we see that approaches increasingly a geodesic ball centered at with radius as This also implies that
-Combining the above, we obtain
-Therefore, there exists a constant such that
-holds for all sufficiently large ∎
-Corollary 4.5 (Estimates of harmonic measures).
-We have
-If then
-If then
-for all sufficiently large In the above, is the Riemannian volume element of
+According to the definition of Green function for it is immediate that for
+Since for then we obtain
+Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [8] proved the following theorem.
+Lemma 4.4.
+Let be a complete Riemannian manifold. Let be a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
+holds for any non-negative harmonic function on where is the Riemannian distance between and and is the lower bound of Ricci curvature of
+Theorem 4.5.
+There exists a constant such that
 Proof.
-The corollary follows immediately from the following relation
-4.2.2. Calculus Lemma
-In order to establish the calculus lemma which plays an important role in the proof of the second main theorem later, we need to estimate the bounds of the Green function for in terms of integral forms.
-Set
-Lemma 4.6 (Estimates of Green functions).
-We have
-If then
-If then for any there exists sufficiently large such that for we have
-Proof.
-Theorem 4.7 (Calculus Lemma).
-Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a positive constant and a subset with finite Lebesgue measure such that
+By
+we obtain an estimate for upper bounds of as follows.
+Corollary 4.6.
+There exists a constant such that
+where is the Riemannian area element of
+4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
+Lemma 4.7 (Borel Lemma).
+Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
 holds for all outside
 Proof.
-If then is the geodesic disc centered at with radius It yields from Theorem 4.6 that
+The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit exists or equals If then Set
+Note that exists for almost all Then, we have
+This completes the proof. ∎
 Set
-Then
-which yields from Corollary 4.5 that
-On the other hand, applying Borel lemma to twice, then for any there exists a subset with finite Lebesgue measure such that
-holds for all outside Combining (15) with (16), we have
-holds for all outside where is a constant.
-Next, we show the case where Invoking Lemma 4.6, for any there exists sufficiently large such that for we have
-Notice that as due to Lemma 4.3, then there exists sufficiently large such that for we have
-Hence, we conclude that
+Theorem 4.8 (Calculus Lemma).
+Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is given by
+Proof.
+Invoking Lemma 4.3, we have
+Set
+A simple computation leads to
+In further, we have
+Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
+holds for all outside On the other hand, Corollary 4.6 implies that there exists a constant such that
+Set Combining the above, we have
+holds for all outside ∎
+In the following, we establish a logarithmic derivative lemma. Let be a meromorphic function on The norm of the gradient of is defined by
+in local holomorphic coordinates here is the inverse of Define the Nevanlinna’s characteristic function of by
 where
-Set
-A direct computation leads to
-Furthermore, we have
-According to Corollary 4.5, there exists sufficiently large such that
-holds for all Thus, it follows from (18) that
-Using Borel lemma twice, for any , we have
-holds for all outside a subset with finite Lebesgue measure. Applying (19) and (20) and noting that as we see that there exists sufficiently large such that for all we have
-Therefore, by this with (4.2.2)
-holds for all outside where
-is a constant. This proves the theorem. ∎
-4.2.3. Logarithmic Derivative Lemma
-To establish the second main theorem, we still need to prove a logarithmic derivative lemma.
-Let be the gradient operator on associated with the Kähler metric Let be a meromorphic function on The norm of the gradient of is defined by
-in a local holomorphic coordinate where is the inverse of Define
-where
-Indeed, define for
-where
-It is not difficult to check that
-On take a singular metric
-which is satisfied with
-To establish the Logarithmic Derivative Lemma, we require the following two lemmas.
+It is not difficult to show that
+On one puts a singular metric
+which leads to
 Lemma 4.9.
@@
 Proof.
-A direct computation leads to
-which yields from Fubini’s theorem that
+Note that
+Whence, it concludes from Fubini’s theorem that
 Lemma 4.10.
-Assume that Then for any there exists a subset with finite Lebesgue measure such that
-holds for all outside
+Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is given by
 Proof.
-Make use of the concavity of , we have
-Thus, it yields from Theorem 4.7 that
-for all outside The lemma is proved. ∎
 Define
 Theorem 4.11 (Logarithmic Derivative Lemma).
-Let be a nonconstant meromorphic function on Then for any there exists a subset with finite Lebesgue measure such that
-holds for all outside
+Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is given by
 Proof.
 Note that
-By Lemma 4.9, for any there exists a subset with finite Lebesgue measure such that
+By Lemma 4.18, we have for any there exists a subset of finite Lebesgue measure such that
 holds for all outside ∎
@@
 Theorem 4.12 (Second Main Theorem).
-Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
-holds for all outside
+Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is given by
 Proof.
@@
 is a singular volume form on By we obtain
-in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1), it yields that
-On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and with On each we write
+in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1)
+On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and
+for On each write
 where is a positive smooth function on Given a partition of the unity subordinate to Put
 Again, set Then on we have
-Fix a we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
+For any we can take local holomorphic coordinates near and a local holomorphic coordinate near such that
 It deduces that
@@
 on Since is bounded on then it yields that
-By this with Dynkin formula (see Lemma 4.1) and Theorem 4.10, for any there is a subset with finite Lebesgue measure such that
-holds for all outside By (4.2.4) and (23), then the theorem holds. ∎
-Let us consider the defect relation. Define the defect and the simple defect of with respect to respectively by
-According to Theorem 4.2, we obtain Since has maximal volume growth, then whenever is nonconstant, as
-Corollary 4.12 (Defect relation).
-Assume the same conditions as in Theorem Then
+By this with Dynkin formula and Theorem 4.11, for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where one used the fact that Combining (4.2.4) and (19), we have the theorem proved. ∎
 4.3. Second Main Theorem for Non-positive Sectional Curvature
-In the subsection, we assume that has non-positive sectional curvature.
-Fix a point Review that denotes the geodesic ball centered at with radius By Sard’s lemma, the boundary of is a submanifold of for almost every Set
-where is the pointwise lower bound of Ricci curvature defined by
-It remains to use to denote the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and use to denote the harmonic measure on with respect to Let be a meromorphic function.
-Similarly, the Nevanlinna’s functions are defined by
-Refer to Dong [14] (see Atsuji [3] also), we have the Second Main Theorem as follows:
+Assume that has non-positive sectional curvature satisfying (3).
+From the definition of curvatures, it is not difficult to see that In particular, we have if has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering As long as is equipped with the pull-back metric induced from via then has the same curvatures as ones of In fact, is a Cartan-Hadamard manifold.
+In the following, we denote by the Riemannian distance function of from on
+4.3.1. Construction of and Nevanlinna’s functions
+A. Setups for
+Due to the curvature conditions, we have a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
+and
+Evidently, the two ODEs have the unique solutions
+Using the estimates of we have (see [17, 34])
+where
+Here, is the standard Euclidean area of For define
+Since as and then we deduce that is a precompact domain containing in which exhausts So, the boundary of can be formulated as
+Set
+which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. Denote by the harmonic measure on with respect to i.e.,
+where is the inward normal derivative on is the Riemannian area element of In particular, if In this case, we have
+which is just the geodesic ball centered at with radius in Then
+Let be the geodesic ball centered at with radius in We give the relation between and as follows.
+Set
+Theorem 4.13.
+For we have and
+Hence, there exists a constant independent of such that
+Proof.
+B. Setups for
+We shall extend the settings for to the case that For the function does not make sense. However, since
+we can extend continuously to defined by (4).
+Notice that is parabolic when i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case.
+However, we note (21) for it is thus natural to define
+for Let it then yields from (22) that
+To conclude, has the unified form
+for In further, we have
+for
+Let be a meromorphic function. Similarly, one can defined the Nevanlinna’s functions and of as well as on as in Section 4.2.1 and Section 4.2.4.
+4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
+Lemma 4.14.
+For we have for
+Proof.
+When is the geodesic ball centered at with radius The conclusion holds immediately due to (23). In the following, we consider the other case. Since on then for
+This completes the proof. ∎
+Theorem 4.15.
+There exists a constant such that
+holds for
+Proof.
+When is a geodesic ball centered at with radius Hence, it yields from (4) and (23) that
+This implies the conclusion holds. For by Lemma 4.4 with (3), (4) and (20) (see Remark 5 in [34] also)
+for some large constant which depends only on the dimension By (20) again
+which gives for So, we conclude that
+where ∎
+Corollary 4.16.
+There exists a constant such that
+holds for where is the Riemannian area element of
+Set
+Theorem 4.17 (Calculus Lemma).
+Let be a locally integrable function on Assume that is locally bounded at Then for any there exists a set of finite Lebesgue measure such that
+holds for outside where is defined by
+Proof.
+Lemma 4.14 yields that
+Differentiating to obtain
+In further
+Applying Borel’s lemma to the left hand side on the above equality twice, then we see that there exists a set of finite Lebesgue measure such that
+holds for outside Meanwhile, Corollary 4.16 means that there exists a constant such that
+Substituting (27) into (4.3.2), we conclude that
+for outside , where ∎
+Lemma 4.18.
+Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
+holds for outside where is defined by
+Proof.
+Theorem 4.19 (Logarithmic Derivative Lemma).
+Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
+holds for outside where is defined by
+Proof.
+4.3.3. Second Main Theorem
 Theorem 4.20 (Second Main Theorem).
-Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
+Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside where is defined by
+Proof.
+Applying the similar arguments as in the proof of Theorem 4.12 and combining Theorem 4.17 with Theorem 4.19, we conclude that for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
-Similarly, we can define the defect and the simple defect of with respect to a value in
-Corollary 4.14 (Defect relation).
-Assume the same conditions as in Theorem If satisfies the growth condition
-then
-5. Nevanlinna theory of algebroid functions on Kähler manifolds
-Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to its Kähler metric We adopt the same notations such as etc. defined in Section 4.
+In the following, we estimate When we have
+This implies that the theorem holds. We consider the other case that Since as we deduce that
+Thus, we conclude that
+This completes the proof. ∎
+5. Nevanlinna Theory of Algebroid Functions
+Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
 5.1. First Main Theorem
-Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definition in Section 4.1).
+Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definitions in Section 4.1).
 In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
-Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus Suppose The characteristic function of for is defined by
+Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus We can choose such that
+The characteristic function of for is defined by
 where
@@
 In order to unify the form, we write
-We show the First Main Theorem:
 Theorem 5.1 (First Main Theorem).
@@
 Proof.
-With the aid of Vieta’s theorem, we obtain
-Using Dykin formula (see Lemma 4.1), then it yields that
-From the definition, we see that the left hand side equals
+With the aid of Vieta’s theorem, we obtain Dynkin formula yields that
+By definition, the left hand side equals
 On the other hand, Theorem 2.22 and Theorem 2.23 imply that the right hand side equals
-Thus, we obtain
+We obtain
 Set then
-where
+with
 In particular, and Similarly, we have
@@
 Since
-and
 then we obtain
@@
 where
-Recall that the discriminant of is given by
-which is a -order determinant, see Section 2.1.
+Review that the discriminant of is given by
 Lemma 5.2.
@@
 Take any Let be the -valued, -valued components of at respectively, where Since the set of singular branching points of is an analytic set of complex dimension not greater than then one just needs to consider that is a non-singular branching point of By Theorem 2.14, we can write for
-in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of One can note from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.23, we have
+in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of So, one can derive from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.22, we have
 This implies the lemma holds. ∎
@@
 5.3. Second Main Theorem
-Let be the -leaf complex manifold of -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.22.
+Let be the -leaf complex manifold of a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.21.
 Let be the natural projection.
-Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on for that could be not positive-definite at points in the branching set of Using Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we show that Dykin formula (see Lemma 4.1) still holds under the pullback metric Set Pick a sequence of neighborhoods of such that
-Equip with a complete Kähler metric Let Clearly, is a complete Kähler metric on and as Let be the Laplace-Beltrami operators of respectively.
-Then, weakly as Take a bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and by the harmonic measures of on with respect to respectively.
-It is evident that as and (weakly) uniformally on as due to
+Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula (see Lemma 4.1) still holds under the pullback metric
+Set Picking a sequence of neighborhoods of such that
+Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then we may assume that (otherwise, a subsequence can be chosen so that as ).
+Let be the Laplace-Beltrami operators of on respectively.
+We have weakly as Take an arbitrary bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and again by the harmonic measures of on with respect to respectively.
+It is not difficult to show that as and (weakly) uniformally on as due to
 for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively.
-To unify the form, one uses to denote the volume elements of for respectively.
-Using Dynkin formula, it gives that
-for a -class function with on outside a polar set of singularities at most. Hence, it yields immediately by letting that
-This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.11 and Theorem 4.13 still hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
+To unify the form, one uses to stand for the volume elements of for respectively.
+Apply Dynkin formula, it gives
+for a -class function with on outside a polar set of singularities at most. Hence, it yields by letting that
+This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
 5.3.1. with non-negative Ricci curvature
-In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth.
-With the pullback metric has non-negative Ricci curvature and maximal volume growth.
-Let be the Kähler form of Take where is a bounded domain defined in Section 4.2.
-Let us recall the Green function for and the harmonic measure on defined in Section 4.2.
+Assume that is non-parabolic.
+With the pullback metric is also non-parabolic, with non-negative Ricci curvature.
+Let be the Kähler form of Take where is a bounded domain defined in Section 4.2.1.
+Recall the Green function for and the harmonic measure on defined in Section 4.2.1.
 Then, the characteristic function of is defined by
@@
 and
-If is not constant, then it follows from Theorem 4.11 that for distinct values there is a subset with finite Lebesgue measure such that
+If is not constant, then it follows from Theorem 4.12 that for distinct values there is a subset of finite Lebesgue measure such that
 holds for and all outside
 Theorem 5.4 (Second Main Theorem).
-Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
+Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
 Proof.
-Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (1), say where with and gives a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
+Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (7), say where with and is a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
 On the other hand, we have and
-Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset with finite Lebesgue measure such that
+Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
 holds for all outside Hence, we have the theorem proved. ∎
-We continue to treat the defect relation. Define the defect and the simple defect of with respect to respectively by
-According to Theorem 5.1, we obtain The condition of maximal volume growth gives that if is nonconstant, as
-Corollary 5.5 (Defect relation).
-Assume the same conditions as in Theorem Then
-Corollary 5.6 (Picard Theorem).
-Let be a nonconstant -valued algebroid function on Then, can omit distinct values at most in
+When we have By it is trivial to deduce Hence, we obtain:
+Corollary 5.5.
+Let be a nonconstant -valued algebroid function on with Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside
 5.3.2. with non-positive sectional curvature
-In the subsection, we assume that has non-positive sectional curvature.
-Then, also has non-positive sectional curvature under the pullback metric Take where is the geodesic ball centered at with radius Recall the Green function for and the harmonic measure on in Section 4.3.
-Similarly, the characteristic function of is defined by
+Assume that has non-positive sectional curvature satisfying (3).
+Note that the curvatures of also satisfy (3) under the pullback metric In the same way, we obtain notations and by setting Apply the similar arguments as in the proof of Theorem 5.4, we can show that
+Theorem 5.6 (Second Main Theorem).
+Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside
+If then and Hence, we conclude that
+Corollary 5.7.
+Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
+holds for all outside
+Let be the unit ball centered at the origin 0 in Consider equipped with the Poincaré metric
+in holomorphic global coordinates where
+It is known that is a non-compact complete Kähler manifold of constant sectional curvature under this metric. From (29), the associated Laplace-Beltrami operator is
+Take as a reference point. The associated Poincaré distance function of from 0 on can be easily written as
+Hence, we have the minimal positive global Green function of for with a pole at 0:
+where denotes the standard Euclidean area of the unit sphere in By definition (see Section 2.1), we have
+This gives
+which is just the geodesic ball centered at 0 with radius in Hence, the boundary of is the geodesic sphere centered at 0 with radius Whence, we obtain the Green function of for with a pole at 0 satisfying Dirichlet boundary condition:
+So, for with
+We write
+in the (geodesic) polar coordinate form, where is the Riemannian metric on induced by and
+is the standard Euclidean metric on It yields from (29) that
 where
-Indeed, for
-where
-In order to unify the form, we write
-Also, define for
-Assume that We denote by the connected components of centered at respectively, here for Again, define
-We see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and is the harmonic measure on with respect to Put
-with
-where is the Ricci curvature tensor on Similarly, set
-where
-Also, define for
-If is not constant, then it follows from Theorem 4.13 that for distinct values there is a subset with finite Lebesgue measure such that
-holds for and all outside
-Apply the similar arguments as in the proof of Theorem 5.4, we obtain:
-Theorem 5.7 (Second Main Theorem).
-Let be a nonconstant -valued algebroid function on Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
-holds for all outside
-Set
-Corollary 5.8 (Defect relation).
-Assume the same conditions as in Theorem If satisfies the growth condition then
-Corollary 5.9 (Picard Theorem).
-Let be a nonconstant -valued algebroid function on satisfying the growth condition then can omit distinct values at most in
-6. Unicity problems of algebroid functions on Kähler manifolds
-In this section, is a non-compact complete Kähler manifold of complex dimension
+Let be the Riemannian area element of Using the second formula in (31) and volume form formula, it gives immediately
+where is the standard Euclidean area element of
+With these preparations, we shall give an estimate for Ricci term Let be the scalar curvature of which is a negative constant. It follows from (32) that
+Substituting (30) into the expression of and calculating the integral, it is not hard to deduce that
+We proceed to estimate with By
+we have Hence, we obtain:
+Corollary 5.8.
+Let be a nonconstant -valued algebroid function on Equip with the Poincaré metric. Then
+holds for all outside a subset of finite Lebesgue measure.
+6. Unicity Problems of Algebroid Functions
+Assume that is a non-compact complete Kähler manifold.
 6.1. Propagation of Algebraic Dependence
-Let be the Fubini-Study form on Given a -valued algebroid function on and Define
-Using Dynkin formula (see Lemma 4.1), it is immediate that
-On the other hand, since
-then we have
+Let be the Fubini-Study form on Given a -valued algebroid function on For define
+It is not hard to deduce that Since and we obtain
 Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable.
@@
 Definition 6.1.
-Let be an analytic subset of nonconstant algebroid functions on are said to be algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, we say that are -related on
+Let be an analytic subset of We say that nonconstant algebroid functions on are algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, are said to be -related on
 Let be an effective divisor on and fix any positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
@@
 Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ).
-We introduce the notation as follows:
-If has non-negative Ricci curvature and maximal volume growth, then we denote by
+We introduce the notation as follows.
+When is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), denote by
 the set of all such that
-If has non-positive sectional curvature, then we denote by
+When has non-positive sectional curvature satisfying (3), denote by
 the set of all such that
-and
-Let stand for the hyperplane line bundle over i.e., is the holomorphic line bundle defined by a point in It gives a holomorphic line bundle over via
+and (5) is satisfied.
+Let be the hyperplane line bundle over It gives a holomorphic line bundle over
 where is the natural projection on the -th factor with Let for For , define
 with
-for where (the first Chern class of ) is a fixed positive -form.
+where is a fixed positive -form.
 Lemma 6.2.
@@
 A direct computation leads to
-Then, it follows that
-Invoking Dynkin formula (see Lemma 4.1), it is therefore
-This proves the lemma. ∎
+Thus, we have
+Apply Dynkin formula, we deduce that
+This completes the proof. ∎
 Let be the set of all indecomposable hypersurfaces in such that for some Put
@@
 Set
-with
-We obtain a criteria for the propagation of algebraic dependence:
+and
+where
 Theorem 6.5.
@@
 Fix any -valued algebroid function We say that is generic with respect to if for at least one Assume that is generic with respect to Let be the set of all such that on Set
-The following is a unicity theorem:
 Theorem 6.6.
@@
 6.2. Five-Value Type Theorem
-We shall apply the propagation theory of algebraic dependence established in Section 6.1 to the unicity problems of algebroid functions on non-compact complete Kähler manifolds under two kinds of different curvature conditions.
-A1 has non-negative Ricci curvature and maximal volume growth
-Under this curvature condition, we show that
+A1 has non-negative Ricci curvature
+Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
 Theorem 6.7.
@@
 Set and with Then
-Making use of Theorem 6.6, we have holds. Similarly, if one sets for then we obtain
-which also implies holds. ∎
+By Theorem 6.6, we have holds. Similarly, when for we deduce that
+which implies holds. ∎
 Corollary 6.8 (Five-Value Type Theorem).
@@
 B1 has non-positive sectional curvature
-Let be defined by (24). Set
-Apply the similar arguments as in the proof of Theorem 6.7, we obtain:
+Assume that has non-positive sectional curvature satisfying (3).
 Theorem 6.9.
-Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy the growth condition Then
+Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that satisfy Then
 Assume for all If then
@@
 Proof.
-The theorem can be confirmed by carrying the same arguments as in the proof of Theorem 6.7. ∎
+The argument is similar as in the proof of Theorem 6.7. ∎
 Corollary 6.10 (Five-Value Type Theorem).
-Let be nonconstant -valued algebroid functions on Assume that satisfy the growth condition If share distinct values ignoring multiplicities in then
-Remark 6.11.
-According to the well-known fact from the case for we see that the number of the shared values is sharp in Corollaries 6.8 and 6.10.
-However, Theorems 6.7 and 6.9 cannot yield the optimal number of the shared values if In doing so, we need to improve Theorems 6.7 and 6.9 so that one can derive the optimal number of the shared values.
-A2 has non-negative Ricci curvature and maximal volume growth
-To a general unicity problem for the value sharing ignoring multiplicities, we establish the following theorem:
+Let be nonconstant -valued algebroid functions on Assume that satisfy If share distinct values ignoring multiplicities in then
+A2 has non-negative Ricci curvature
+Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
 Theorem 6.11.
-Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies
-with then
+Let be nonconstant -valued, -valued algebroid functions on respectively. Let be distinct values in Assume that for all If satisfies then
 Proof.
@@
 in the sense of currents, where and are understood as divisors. So, we are led to
-where It yields from Lemma 6.3 that
+It yields from Lemma 6.3 that
 Similarly,
-Put together (6.2) and (6.2), it yields from Theorem 5.4 that
-Therefore,
-But, it contradicts with the assumption of The theorem is proved. ∎
+Combining (6.2) with (6.2), it yields from Theorem 5.4 that
+It is therefore
+But, it contradicts with the assumption for This completes the proof. ∎
 Corollary 6.12.
@@
 Proof.
-Corollary 6.13 leads to a sharp five-value type theorem:
 Corollary 6.13 (Five-Value Type Theorem).
@@
 B2 has non-positive sectional curvature
-Using the similar arguments as in the proof of Theorem 6.12, we obtain:
+Assume that has non-positive sectional curvature satisfying (3).
 Theorem 6.14.
-Let be nonconstant -valued, -valued algebroid functions on respectively satisfying the growth condition Let be distinct values in Assume that for all If satisfies
-with then
+Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Assume that for all If satisfies then
+Proof.
+The argument is similar as in the proof of Theorem 6.11. ∎
 Corollary 6.15.
-Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition Let be distinct values in Then
+Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. Let be distinct values in Then
 Assume for all If then
 Assume for all If then
-Corollary 6.16 leads to a sharp five-value type theorem:
 Corollary 6.16 (Five-Value Type Theorem).
-Let be nonconstant -valued, -valued algebroid functions on respectively, satisfying the growth condition If share distinct values ignoring multiplicities in then
+Let be nonconstant -valued, -valued algebroid functions on satisfying respectively. If share distinct values ignoring multiplicities in then

### v10

@@
 Abstract.
-We generalize the classical Nevanlinna theory for algebroid functions on to a complete Kähler manifold with non-negative Ricci or non-positive sectional curvature.
-As an application, we give some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
+In this paper, we generalize the classical Nevanlinna theory of algebroid functions from to a complete Kähler manifold with either non-negative Ricci curvature or non-positive sectional curvature.
+As its applications, we establish some Picard type theorems and five-value type theorems for algebroid functions under certain conditions.
 Key words and phrases:
@@
 The notion of algebroid functions (see [16, 28]) introduced by H.
-Poincaré is a generalization of algebraic functions and meromorphic functions.
-Let us review that a -valued algebroid function on a (nonempty) domain (the complex plane) is defined by an irreducible algebraic equation
-where are holomorphic functions with no common zeros on An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G.
+Poincaré is an extension of meromorphic functions.
+Recall that a -valued algebroid function on a domain is defined by an irreducible algebraic equation
+where are holomorphic functions without common zeros in An algebroid function is viewed as a special multi-valued analytic function, while which was first put forward, G.
 Darboux deemed that it is an important class of functions.
-To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a (generic) -sheeted analytic covering of A further fact claims that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
+To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a generic -sheeted analytic covering of A further fact states that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
 The most concerned problem on algebroid functions seems to be the value distribution.
@@
 Rémoundos [24] proved a Picard’s theorem which claims that a nonconstant -valued algebroid function on omits distinct values at most in Afterwards, G.
-Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to this topic, including investigations to Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc.
-More details refer to H.
+Valiron [40] showed without details a five-value type theorem in unicity theory, which claims that two nonconstant -valued algebroid functions on will be identical if they share +1 distinct values ignoring multiplicities in Since then, a large number of scholars have devoted themselves to the study of Nevanlinna theory, Ahlfors theory of covering surfaces, unicity theory and normal family theory, etc.
+for algebroid functions.
+We refer readers to H.
 Cartan [7], He-Xiao [16], Niino-Ozawa [19], J.
@@
 Algebroid function theory in one complex variable didn’t mature until the last 30-40 years.
-However, there is little literature on the algebroid functions in several complex variables, except for the work of Hu-Yang [15].
-Given two nonconstant algebroid functions on a non-compact Hermitian (Kähler) manifold .
-We are inspired by the following three fundamental problems:
+Up to now, there seems to be very little literature regarding algebroid functions in several complex variables, except for the limited work by Hu-Yang [15].
+Let denote any two non-constant algebroid functions on a non-compact Hermitian or Kähler manifold We are inspired by the following three fundamental problems:
 How to realize the uniformization of ?
@@
 How many values should share at least to guarantee ?
-In this paper, we answer these three problems by generalizing the classical Nevanlinna theory for algebroid functions on to a non-compact complete Kähler manifold with non-negative Ricci or non-positive sectional curvature.
+In the present paper, we answer these problems. What is more important is that we develop the classical Nevanlinna theory of algebroid functions on to complete non-compact Kähler manifolds.
 1.2. Main Results
 A. Uniformization
-Let be a non-compact Hermitian manifold of complex dimension At first, we generalize the notion of algebroid functions from to by means of a technical treatment (see Definition 2.15).
-Let be a -valued algebroid function on Assume that is a -leaf meromorphic function element of namely, is a local -valued meromorphic solution of equation (7) in a neighborhood of where is a branching point with order of If is a “non-singular” branching point of then we describe the asymptotic behavior of as approaches i.e., we prove that can be expanded into the Pusieux series (see Theorem 2.14)
-with around in local holomorphic coordinates where are meromorphic functions in With this Pusieux series, we show that
+Let be a non-compact Hermitian manifold.
+We generalize the notion of algebroid functions from to (see Definition 2.15).
+Let be a -valued algebroid function on Given a -leaf function element of When is not a singular branching point of we give a local Pusieux expansion of at i.e., in a local holomorphic coordinate we show that can be expanded locally into the following convergent Pusieux series (see Theorem 2.14)
+at with where are meromorphic functions. By this Pusieux series, we construct a -leaf complex manifold and prove that
 Theorem A.
-Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
+Let be the -leaf complex manifold by a -valued algebroid function on Then, can lift to a meromorphic function on via the natural projection such that
 B. Nevanlinna theory
-In this part and the following Parts C, D, we utilize to stand for a non-compact complete Kähler manifold.
-Let be a -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed reference point To unify the form, we write
+In this part and the following Parts C, D, we use to stand for a complete non-compact Kähler manifold.
+Let be any -valued algebroid function on Refer to the definition in Section 5.3, we have the Nevanlinna’s functions of and on where is the Ricci form of and is a precompact domain containing a fixed point To unify the form, we write
 The simple defect of with respect to is defined by
@@
 is of non-polynomial type growth, i.e., satisfies the growth condition
-We have as due to
+Note that as due to
 The condition (2) is relaxed. For example, it holds when with and holds when with or and
@@
 holds for all outside where is defined by
-Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
+Note that Theorem B4 covers the classical second main theorem for both and the unit ball with Poincaré metric in (see Corollaries 5.7 and 5.8).
 Corollary B5 (Defect Relation).
@@
 and (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has non-positive sectional curvature satisfying (3).
-Let be the hyperplane line over We denote by the set of all indecomposable hypersurfaces in such that for some where and is the natural projection on the -th factor. Put
+Let be the point line bundle over We denote by the set of all indecomposable hypersurfaces of such that for some in which and is the natural projection on the -th factor. Set
 with and
@@
 First of all, we would like to extend the notion of algebroid functions to a non-compact Hermitian manifold.
-Let be a -dimensional non-compact Hermitian manifold.
+Let be an -dimensional non-compact Hermitian manifold.
 Consider an irreducible algebraic equation in on :
 where are holomorphic functions locally defined on such that they well define a meromorphic mapping
-Let be the indeterminacy locus of which is an analytic set in of complex codimension not less than if or an empty set if So, is a holomorphic mapping of into
+Let be the indeterminacy locus of which is either an analytic set of complex codimension not smaller than for or an empty set for in Then, is holomorphic on
 Remark 2.1.
 The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for all satisfying that
-where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on for
+where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on with
 2.1.1. Meromorphic function elements
@@
 Definition 2.2.
-Let A pair is called a regular function element, if is holomorphic in a neighborhood of A pair is called a meromorphic function element, if is meromorphic in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
+Let be any point in A pair is called a regular function element, if is a holomorphic function in a neighborhood of and called a meromorphic function element, if is a meromorphic function in a neighborhood of A pair is said to be subordinate to equation if is a local solution of equation in a neighborhood of
 Theorem 2.3.
 Let be a pair such that
-Then there exists a unique regular function element subordinate to equation satisfying
+Then, there exists a unique regular function element subordinate to equation satisfying
 Proof.
-The theorem can be showed by employing implicit function theorem. Here, we present an alternative proof of complex analysis. Taking a such that (see Remark 2.1). We write in the form
+The theorem can be confirmed by applying implicit function theorem. Here, we give an alternative proof in complex analysis. Take a such that given by Remark 2.1. Write as the form
 where
-The conditions mean that
-By the continuation, there exist such that
-for all satisfying and In addition, those are in In further, restrict to then we obtain
-Employing Rouché’s theorem, for any fixed with and have the same number of zeros in the disc In fact, the number of zeros is 1, since in has only one zero in the disc when This implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that
-Evidently, is continuous on
-Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking local holomorphic coordinates around such that Apply Hartog’s theorem, it is sufficient to show that is holomorphic in every holomorphic direction at For and , set
+By the conditions, we have
+The continuity implies that here exist such that
+for all with and In further, restrict to then we obtain
+According to Rouché’s theorem, we deduce that for all with and have the same number of zeros in the disc In fact, this number of zeros is 1, since in has the unique zero in the disc for all with It implies that uniquely determines a function in a smaller geodesic ball neighborhood of with radius such that Evidently, is continuous on
+Next, we show that is holomorphic on One just needs to show without loss of generality that is analytic at Taking a local holomorphic coordinate near such that Thanks to Hartog’s theorem, it is sufficient to examine that is holomorphic in each holomorphic direction at For and put
 Note that
@@
 For any the fundamental theorem of algebra asserts that
-has exactly distinct complex roots. Therefore, Theorem 2.3 implies that
+has exactly distinct complex roots. Hence, Theorem 2.3 implies that
 Corollary 2.5.
@@
 To see the common roots of and we define the resultant of and by a -order determinant:
-where for According to Sylvester’s elimination method, we can see that if and only if or has a multiple root as a polynomial in i.e., This gives that is an analytic set of complex dimension Define the discriminant of by
+where for Apply Sylvester’s elimination method, we can deduce that if and only if or has a multiple root as a polynomial in i.e., Thus, it means that is an analytic set of complex codimension Define the discriminant of by
 Assume that splits into linear factors, say
@@
 If then has a multiple root as a polynomial in If we set and rewrite (7) as where
-Since then one can see that is a zero of i.e., a pole of In fact, we will deduce from Theorem 2.22 (in Section 2.3) that if and only if To conclude, we see that Corollary 2.5 can be extended to meromorphic function elements as follows:
+Since we see that is a zero of i.e., a pole of In fact, we note from Theorem 2.21 (see Section 2.3) that if and only if To conclude, Corollary 2.5 can be extended to as follows.
 Corollary 2.6.
@@
 2.1.2. Analytic continuation of meromorphic function elements
-Recall that a meromorphic function element on is a pair where is a meromorphic function defined in a neighborhood of in As a matter of convenience, we write
-for short without confusion.
-Instead of we sometimes take a geodesic ball neighborhood centered at with radius Let be any two meromorphic function elements on Define an equivalent relation as follows: if and only if
+We recall that a meromorphic function element on is a pair such that is a meromorphic function defined in a neighborhood of For convenience, we write without any confusion.
+We usually take which is a geodesic ball neighborhood centered at with radius For two meromorphic function elements and define an equivalent relation: if and only if
 Set
 We equip with a topology.
-A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we can regard that
-A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in Namely, In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying that In other words, it is defined to be A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It is not hard to check that it defines a topology of
+A meromorphic function element is said to be a direct analytic continuation of a meromorphic function element if and in a neighborhood of By the analytic continuation, we may regard that
+A neighborhood of is defined by a set (containing ) of all such that is a direct analytic continuation of satisfying where is a neighborhood of in That is,
+In particular, a -neighborhood of (satisfying that ), is defined by the set (which contains ) of all such that is a direct analytic continuation of satisfying In other words, it is defined by A subset is called an open set, if either is an empty set or every point is an inner point, i.e., there is a -neighborhood of such that It is not hard to check that it defines a topology of
 Theorem 2.7.
@@
 Proof.
-Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element It implies that in a small neighborhood of By the uniqueness theorem of analytic functions, we obtain but it contradicts with ∎
-Next, we introduce the notion of analytic continuation (by which we mean both holomorphic continuation and meromorphic continuation) of meromorphic function elements.
+Pick If then we can consider -neighborhoods of and of where Evidently, we have If for then we can take a function element which gives that in a smaller neighborhood of Using the uniqueness theorem of analytic functions, we obtain a contradiction with ∎
 Definition 2.8.
@@
 Remark 2.9.
-The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve.
-Thus, the analytic continuation is symmetric.
+The analytic continuation of elements in along a curve in is a continuous mapping of into namely, an analytic continuation of to along is a curve connecting and Moreover, if is an analytic continuation of along a curve, then is also an analytic continuation of along the same curve, i.e., the analytic continuation is symmetric.
 Theorem 2.10 (Analytic continuation).
@@
 Set
-Then, we have We show Otherwise, we assume that By Corollary 2.6, there are exactly function elements at say that
-Take small enough such that where Then there is a unique integer such that Notice that the domain of definition of is so for any one can redefine function elements Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Thus, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic (meromorphic) functions.
+Evidently, It is sufficient to prove that Otherwise, we may assume that By Corollary 2.6, there exist exactly function elements at denoted by
+Take small enough such that with Then, there is a unique integer such that Notice that the domain of definition of is for each one can redefine function element Combined with the definition of it yields immediately that can continue analytically to along i.e., is an analytic continuation of along However, it contradicts with the definition of Hence, we obtain The uniqueness of is immediate due to the uniqueness theorem of analytic functions.
 ∎
 Corollary 2.11.
-Let be distinct function elements at respectively. Then can continue analytically to each other along any curve in
+Let be any two groups of distinct function elements at respectively. Then, can continue analytically to each other along any curve in
 Theorem 2.12 (Connectivity).
@@
 Proof.
-Let be distinct function elements at every fixed With the aid of Corollary 2.11, then we only need to prove that all can continue analytically to each other.
-Let be the largest integer such that there are function elements in which can continue analytically to each other along a curve.
-Clearly, we have Without loss of generality, we shall assume that these function elements are So, we can write
-where are holomorphic functions in a neighborhood of By Vieta’s theorem, we obtain
-It is equivalent to show that: can continue analytically to the whole Apply Theorem 2.10, can continue analytically to along a curve at any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., continue analytically to Hence, holds.
-If then we are able to determine another set via the rest function elements However, it implies that equation (7) is reducible, which is a contradiction.
+Let be distinct function elements at every point With the help of Corollary 2.11, it is sufficient to prove that all can be continued analytically to each other.
+Let be the largest integer such that there are function elements in which continue analytically to each other along a curve.
+Note that Without loss of generality, one may assume that the function elements are just Then, it can be written as
+where are holomorphic functions in a neighborhood of Using Vieta’s theorem, we obtain
+It suffices to show that
+can continue analytically to the whole
+Apply Theorem 2.10, can continue analytically to along a curve for any where is a permutation of By Vieta’s theorem, determine the same in a small neighborhood of i.e., can continue analytically to Hence, holds.
+If then one can determine another set using the rest function elements However, it leads to that equation (7) is reducible, which is a contradiction.
+This completes the proof.
 ∎
@@
 is a path-connected space.
-2.1.3. Properties of meromorphic function elements near multiple points
-We discuss the local properties of function elements near multiple points. Let be a pair satisfying that and
-Taking a sufficiently small neighborhood of such that on By the continuation of there exists a number such that on Let then there exist distinct regular function elements and in respectively.
-By the use of Theorem 2.10, then for a curve with and are able to continue analytically to along In other words, there exists a permutation of such that can continue analytically to along respectively.
-By symmetrical property, can also continue analytically to along However, will not, in general, continue analytically back to along respectively.
-In the following, we discuss the analytic continuation of a function element along a closed Jordan curve surrounding a multiple point.
-Take Pick as a closed Jordan curve passing by around Without loss of generality, we only extend analytically along starting from When returns to from along one will attain to a function If then one continues to extend analytically along Repeating this action, we see finally that cycles on and on periodically.
-Let be the smallest period such that returns to itself, we obtain We say that is a branching point with order (multiplicity) of if Repeatedly, then it is not hard to verify that can be divided into several groups.
-Let be the number of groups.
-If the -th group has members, then we have
-Without loss of generality, we consider the first group: Take a proper we have a biholomorphic mapping such that
-where stands for the polydisc centered at with polyradius in and is a local holomorphic coordinate of in So, we have
-Set and write Next, we shall show that is a single-valued analytic function on where is the unit disc centered at in the -plane.
-Evidently, is holomorphic on We still need to prove that is a removed singularity of In doing so, it is sufficient to show that is bounded on due to Riemann’s analytic continuation.
-There exists a number such that on for Combining (7) with that on we obtain
-on If we obtain Thus, on and this bound does not depend on Whence, is bounded on This means that is holomorphic on and of the Taylor expansion
-near Let be the set of all branching points of in Note that the value of is independent of if is sufficiently small.
-If is a non-singular branching point of i.e., does not lie in a crossing of components of then is smooth whenever is small enough.
-By and Riemann’s analytic continuation (because the boundedness of on and is properly chosen if necessary), there is a biholomorphic mapping with
-such that can be expanded into the Pusieux series
-near where are holomorphic functions in satisfying that for
-According to the above arguments, we show that is a -valued analytic function near In other words, if we cut along a (real) hypersurface which passes through then is separated into single-valued analytic components, and these components replace each other when they continue analytically along a closed curve around in By a local expansion of is a -multiple -valued branching point (with order ) of
-Assume that is not a singular branching point of Set and write
-Similarly, we have
-If then
-Choose such that Put then
-Thus, i.e., the case turns to the case Note that is a pole with order of
-To conclude, for each we have confirmed that there are distinct (single-valued) meromorphic function elements subordinate to equation (7), if is not a branching point.
-If is a branching point, then one shows that there are several multi-valued function elements (which contain single-valued function elements) subordinate to equation (7), say where is a -order branching point of satisfying At this time, we call a -order branching point of If is a non-singular branching point of then we have the convergent Pusieux expansion
-near in a local holomorphic coordinate Note that is a -multiple root of namely, can separate into single-valued components with value at If then can separate into single-valued components which take at In particular, is a meromorphic function element for Furthermore, is a regular function element if also.
-By the analytic continuation, equation (7) uniquely defines a -valued meromorphic function on outside a set of branching points of and is separated into single-valued meromorphic components on Let denote the set of all branching points of which is an analytic set contained in called the branching set of As a result, we conclude that
+2.1.3. Properties of function elements near a multiple point
+Let be a pair such that and
+Let us take a small neighborhood of such that on Using the continuity of for there exists a number such that
+According to Corollary 2.6, for any with there exist two groups of distinct meromorphic function elements at subordinate to equation (7), respectively.
+Apply Theorem 2.10, then for a curve with and we are able to extend analytically to along In other words, there is a permutation of such that can continue analytically to along respectively.
+The symmetry implies that can also continue analytically to along However, we should mention that won’t be, in general, continued analytically back to along respectively.
+In what follows, we would consider the analytic continuation of a function element along a closed Jordan curve surrounding any multiple point Take a smooth closed Jordan curve around such that which is close to and doesn’t surround or cross any connected components of that do not pass through Without loss of generality, we look at the analytic continuation of along starting from When returns to along it will turn to certain function If then one continues to extend analytically along in the same direction of rotation.
+Repeating this action, we shall see that cycles on and on periodically.
+Let be the smallest period such that can return to itself along all such curves Then, we get We call a branching point with order of for Repeatedly, it can be seen that is divided into finitely many groups.
+Let denote the number of groups.
+If the -th group has members, then
+We just analyze the first group: Taking a suitable there exists a biholomorphic mapping such that
+where is the polydisc centered at 0 with polyradius in Setting where is the unit disc centered at 0 in the -plane. For any we have Set and write
+We show that is a single-valued analytic function on Notice that is the branching order of we see that is holomorphic on because is so close to 0 that there exist no any branching points of in We only need to examine that is a removed singularity of To do so, it suffices to prove that is bounded near 0 due to Riemann’s analytic continuation.
+In fact, it follows from (7) and (9) that
+on If we obtain Thus, on i.e., is bounded on It means that is holomorphic on and it has the Taylor expansion
+at Let be the set of all branching points of in Say that a branching point of is non-singular, if does not lie in a crossing of components of and singular otherwise.
+Assume that is not a singular branching point of Then, is smooth if is small enough.
+Taking a suitable if necessary.
+Since is bounded on and defines the branching order of at which is independent of the choice of one can deduce from and Riemann’s analytic continuation that there exists a biholomorphic mapping with
+such that can be expanded locally into the following Pusieux series
+at where are holomorphic functions in Comparing coefficients, we deduce that
+According to the above arguments, we show that is a -valued analytic function defined in a neighborhood of We call a -leaf function element.
+If one cuts along a real hypersurface passing through then can separate into single-valued analytic components, and all these components can turn to each other when they continue analytically along a closed curve around in that is sufficiently close to By the local expansion of we deduce that is a -multiple, -valued branching point with order of
+Assume that is not a singular branching point of Set and define
+By the previous arguments, one can also expand locally into a Pusieux series of the form
+at If then can be expanded locally at in the form
+We take a number with Set Then
+It is clear that
+Hence, the case turns to the case
+To conclude, for each there are distinct multi-leaf function elements subordinate to equation (7), denoted by where is a -valued meromorphic function with and Hence, has branching order at with When we say that is a branching point with order of Moreover, if is not a singular branching point of then we deduce that can expand locally into the convergent Pusieux series at That is, we have shown that
 Theorem 2.14.
-Let be a -valued function element subordinate to equation If is not a singular branching point of then can be expanded into the convergent Pusieux series
-in a local holomorphic coordinate near such that =0 and for where is the branching set in a local neighborhood of and are meromorphic functions in
-We conclude the definition of algebroid functions on as follows.
+Let be a -leaf function element subordinate to equation If is not a singular branching point of then can be expanded locally into the following convergent Pusieux series
+at in a local holomorphic coordinate such that and for all where is the branching set of and are meromorphic functions.
+Note that in Theorem 2.14 is a -multiple root of that is, can separate into distinct single-valued components with value at If then separates into distinct single-valued components which take at
+According to the analytic continuation and uniqueness theorem, the above arguments show that equation (7) uniquely defines a -valued meromorphic function on outside the branching points of Let denote the set of all branching points of which is an analytic set in called the branching set of We generalize the notion of algebroid functions as follows.
 Definition 2.15.
 A -valued algebroid function on is defined by an irreducible algebraic equation
-where are holomorphic functions locally defined on such that they well define a meromorphic mapping from into
-Let denote distinct (single-valued) meromorphic components of Given any -leaf function element of with If denote the single-valued components of then there exists a permutation of such that on In other words, defines a -leaf function element on Thus, any function element of can be represented locally by one or several single-valued components of
-2.2. Multi-leaf Complex Manifolds of Algebroid Functions
-Let be a function element.
-A neighborhood of is defined as a set of all such that is a direct analytic continuation of satisfying i.e., Taking a proper if necessary, there exists a biholomorphic mapping such that where is the unit ball centered at the coordinate origin in Define a natural mapping
-via
-Evidently, is a homeomorphism. Its inverse is
-Set We show that is a complex atlas of Let It suffices to show the transition mapping
+where are holomorphic functions locally defined on such that they define a meromorphic mapping
+In a word, denote by the distinct single-valued meromorphic components of and given a -leaf function element subordinate to equation (7), with distinct single-valued components Then, there is a permutation of such that on
+In the end, we provide an example to illustrate how to calculate the order of branching points of algebroid functions.
+Example. Calculating the order of the -valued algebroid function
+on at branching points
+Since have single-valued components, respectively, we infer that has single-valued components; Since have single-valued components, respectively, we infer that has single-valued components.
+Thus, is a -valued algebroid function.
+Let stand for any closed Jordan curve in All the branching sets of are as follows
+It is clear that
+Order of
+Take such that is surrounded by Note that are the only two branching sets of that pass through Therefore, is not allowed to surround or pass through any other branching sets of except for in order to determine the order of For all such we for
+Thus, the order of is determined by the order of branching point of and the order of branching point of It is
+Order of
+Take such that is surrounded by Note that are the only two branching sets of that pass through Similarly, cannot surround or pass through any other branching sets of except for It implies that for
+Thus, the order of is determined by the order of branching points of It is
+Order of
+Take such that is surrounded by Since are the only three branching sets of that pass through we have for
+Thus, the order of is determined by the order of branching points of respectively. It is
+Order of
+Take such that is surrounded by Since are the only three branching sets of that passes through we have for
+Thus, the order of is determined by the order of branching points of It is
+2.2. Multi-leaf Complex Manifolds by Algebroid Functions
+Let A neighborhood of is defined to be a set of all with such that is a direct analytic continuation of i.e.,
+Take a suitable we have a biholomorphic mapping such that where is the unit ball centered at the coordinate origin 0 in Dfine a natural mapping by
+Clearly, is a homeomorphism whose inverse is
+Set
+We prove that is a complex atlas of Let It suffices to examine that the transition mapping
 is a biholomorphic mapping. This is automatically a homeomorphism. Note that for each we have in a small neighborhood of i.e., A direct computation gives that
-for Similarly, we obtain
-for Hence, is a biholomorphic mapping. On the other hand, for any there corresponds to distinct function elements Therefore, we conclude that
+for any Similarly, we have
+It shows that is a biholomorphic mapping. On the other hand, each corresponds to distinct function elements That is, we conclude that
 Theorem 2.16.
@@
 is a non-compact -leaf complex manifold of complex dimension which is a submanifold of
-In what follows, we shall construct a -leaf complex manifold determined by all functions elements subordinate to equation (7).
+In what follows, we would construct a -leaf complex manifold determined by all function elements subordinate to equation (7).
 Let be any two function elements subordinate to equation (7).
 We define an equivalent relation as follows: if and only if
-for some two single-valued components of respectively, where is the branching set of for Set
-Remark 2.18.
-If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements.
-If then it can correspond to multi-leaf function elements in where is a -order branching point of satisfying In other words, can be separated into single-valued components for
+for some two single-valued components of respectively, where is the branching set of with Set
 Definition 2.18.
-A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one can regard that
+A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one may regard that
 Let A neighborhood of is a set which contains the following elements:
-all such that is a direct analytic continuation of satisfying
-all satisfying
-where is a neighborhood of In particular, a -neighborhood of (satisfying ), is defined by the set which contains the following elements:
-all such that is a direct analytic continuation of satisfying
-all satisfying
-A subset is called an open set, if either is an empty set or every point is an inner point, namely, there exists a -neighborhood of such that If is a -leaf function element, then has leaves.
-In this case, we call a -leaf neighborhood of This gives a topology of and this topology contains the topology of
+all with satisfying that is a direct analytic continuation of
+all with Here and before, is an arbitrary neighborhood of
+In particular, a -neighborhood of with is defined by the set that contains the following elements:
+all with satisfying that is a direct analytic continuation of
+all with
+A subset is called an open set, if is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It gives a topology of which contains the topology of If is a -leaf function element, then has leaves at We call a -leaf neighborhood of
 Theorem 2.19.
@@
 Proof.
-Hausdorff property of can be showed using the similar arguments as in the proof of Theorem 4.3.
-Next, we prove that is connected.
-If not, one can write where are two disjoint open subsets of Set It is evident that Notice that are two disjoint open subsets of since is an analytic set of complex codimension This means that is not connected, but it is a contradiction.
+The Hausdorff property of is showed using the similar arguments as in the proof of Theorem 2.7.
+We assert that is connected.
+Otherwise, one may assume that where are two disjoint open subsets of Set It is evident that Since is an analytic set of complex codimension we see that are two disjoint open subsets of This implies that is not connected, which is a contradiction.
 ∎
-Next, let us continue to endow with a complex structure such that is a complex manifold.
-Take any -leaf function element Assume that is a non-singular branching point of Apply Theorem 2.14 ( is properly chosen if necessary), there exists a biholomorphic mapping such that can be expanded into the Pusieux series
-satisfying that and for
+Next, we will equip with a complex structure such that is a complex manifold.
+Taking a -leaf function element Assume that is not a singular branching point of and is chosen suitably.
+Thanks to Theorem 2.14, there exists a biholomorphic mapping such that can be expanded into the Pusieux series
+on with and for all
 Let be a -leaf neighborhood of containing the following elements:
-all such that is a direct analytic continuation of satisfying
-all satisfying
-Let be single-valued connected components of Define a mapping by
-where Note that is a homeomorphism and its inverse is
-where is a positive integer such that Let be the set of all function elements such that is a singular branching point of some multi-valued component of and let be the set of all corresponding singular branching points.
-It is evident that is an analytic set of complex dimension not greater than Set
-We will show that is a complex atlas of Suppose It is sufficient to prove the transition mapping
+all with satisfying that is a direct analytic continuation of
+all with
+Let be distinct single-valued components of Define a mapping by
+for , where We note that is a homeomorphism and its inverse is
+where is an integer such that
+Let be the set of all function elements satisfying that is a singular branching point of Moreover, denotes the set of all the corresponding singular branching points.
+It is clear that is an analytic set of complex codimension not smaller than Set
+We now show that is a complex atlas of Assume that It is sufficient to verify that the transition mapping
 is a biholomorphic mapping.
-This is automatically a homeomorphism.
-Via a similar discussion, one only needs to deal with the situation where is a -leaf meromorphic function element and is a regular function element.
-For every there exists a single-valued component of such that in some neighborhood of i.e., It is clear that because all function elements in are meromorphic.
+It is automatically a homeomorphism.
+Through a simple analysis, we only need to deal with the situation in which is a -leaf meromorphic function element and is a regular function element.
+For any there exists a single-valued component of satisfying that in some neighborhood of i.e., It is clear that since all function elements in are meromorphic.
 A direct computation leads to
-where is a positive integer such that Hence, is a holomorphic mapping. On the other hand, its inverse is
+where is an integer such that
+Thus, is a holomorphic mapping. On the other hand, its inverse is
 where is the -th single-valued component of Evidently, is also a holomorphic mapping. Hence, is a non-compact -leaf complex manifold of complex dimension
-Finally, we treat the singular part Let be the natural projection by mapping a function element into Then, we note that is a generic -sheeted analytic covering such that
-Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via To conclude, we obtain:
+Treat the singular part Let be the natural projection by mapping any function element to Note that is a generic -sheeted analytic covering such that
+Since is an analytic set of complex codimension not smaller than the complex structure of can be viewed as a natural extension of the complex structure of This implies that one can naturally extend the complex structure of to via Hence, we conclude that
 Theorem 2.20.
-is a non-compact -leaf complex manifold of complex dimension which is called the complex manifold of
+is a non-compact -leaf complex manifold of complex dimension We call the complex manifold by
 2.3. Zero and Pole Divisors of Algebroid Functions
@@
 Proof.
-Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors. Thus, we assume that It yields that
+Let be a zero of Since in the sense of divisors, we may assume that It follows that
 where denotes the number of -valued points of a function at Without loss of generality, suppose that are all single-valued components such that which take at and suppose that are all single-valued components such that which take at Then, we have By Vieta’s theorem, it yields that
-Combining this with (10), we see that reaches the maximal order of a pole of Since then we obtain
-which follows that ∎
+Combining this with (11), we see that reaches the maximal order of a pole of Since we obtain
+which means that ∎
+Similarly, we can show by setting that
 Theorem 2.22.
 We have
-Proof.
-Set Apply the similar arguments as in the proof of Theorem 2.22, we can prove the theorem. ∎
 3. Uniformization of Algebroid Functions
@@
 As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold.
-Our main idea comes from the excellent work of W.
-Stoll [35] on algebroid reduction.
+The main idea originates from W.
+Stoll [35].
 3.1. Algebroid Reduction
-Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, namely, it is a proper and surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex dimension It is evident that
-is a proper, surjective, local biholomorphic mapping of sheet number and hence a -sheeted analytic covering. By Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
-is an analytic set satisfying that and Moreover, and for Let the mapping degree of at is defined in such manner: if is a relatively compact neighborhood of such that then
+Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, by which one means a proper, surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex codimension It is clear that
+is a proper, surjective, local biholomorphic mapping of sheet number and thus a -sheeted analytic covering. By R. Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
+is an analytic set satisfying that and Moreover, and for Let The mapping degree of at is defined in such manner: if is a neighborhood of such that then
 which is independent of the choice of Note that and if and only if By W. Stoll [35], for any we have
-Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection mapping. Let be the -fold symmetric tensor product of i.e.,
+Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection. Let be the -fold symmetric tensor product of i.e.,
 where
-for any is called the symmetric tensor product of and is the -order symmetric group. It is evident that is a subspace of -fold tensor product space with Again, put
+is called the symmetric tensor product of and stands for the -order symmetric group. Note that is a subspace of -fold tensor product space with Again, put
 where for any Naturally, is embedded into by the identity mapping Thus, we obtain
@@
 Definition 3.3.
-Let be a meromorphic mapping with indeterminacy set Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
+Let be a meromorphic mapping with indeterminacy locus Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
 Theorem 3.4 (Stoll, [35]).
@@
 which yields that If then (13) defines an algebraic equation
-Let be the maximal integer in such that Then, equation (13) reduces to Let be the roots of the equation. Apply Vieta’s theorem, it is immediate that
+Let be the maximal integer in such that Then, equation (14) reduces to Let be the roots of the equation. Applying Vieta’s theorem, it is immediate that
 So, is decomposable, i.e.,
-By this with (11), we have To conclude, we obtain
+By this with (12), we have Hence, we obtain
 Lemma 3.6.
@@
 Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function which separates the fibers of such that is the algebroid reduction of via
-Theorem 3.7 (Uniformization theorem).
+Theorem 3.7 (Uniformization).
 Let be a -valued algebroid function on Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
 Proof.
-Let be defined by equation (7). It gives a differentiably non-degenerate and irreducible meromorphic mapping
+Let be defined by equation (7). It gives a differentiably non-degenerate, irreducible meromorphic mapping
 Set with a basis By (15), we can write
@@
 Let be -valued points of Then
-By
-we have the theorem proved. ∎
-Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of generic sheet number via which is a proper surjective holomorphic mapping.
+Combining with then we prove the theorem. ∎
+In conclusion, given a -leaf complex manifold by a -valued algebroid function on there exists the natural projection of generic sheet number given by which is a proper surjective holomorphic mapping.
+That is, we have shown that
 Corollary 3.8.
-Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
+Let be the -leaf complex manifold by a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
 4. Nevanlinna Theory of Meromorphic Functions
-Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
-in local holomorphic coordinates Moreover, we define the Ricci form of by
+Let be a complete non-compact Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
+in a local holomorphic coordinate Moreover, the Ricci form of is defined by
 4.1. First Main Theorem
-Let be an arbitrary relatively compact domain with smooth boundary in Fix a point and denote by the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
+Let be any precompact domain with smooth boundary in Fix a point and denote by the Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
 where is the Dirac’s delta function on with a pole at The harmonic measure on with respect to is defined by
 where is the Riemannian volume element of and is the inward normal derivative on
-Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
+Next, we introduce Nevanlinna’s functions. Let be a meromorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
 The characteristic function of for is defined by
@@
 Proof.
-Apply Dynkin formula (see Lemma 4.1) to we obtain
+Applying Dynkin formula to we obtain
 Note that the first term on the left hand side equals and the right hand side equals
@@
 Assume that is non-parabolic with non-negative Ricci Curvature.
-Fix a reference point and denote by the geodesic ball centered at with radius in Denote by the Riemannian volume of The non-parabolicity of means that
+Fix a reference point and denote by the geodesic ball centered at with radius in Let denote the Riemannian volume of The non-parabolicity of means that
 Thus, we have the unique minimal positive global Green function of for which can be written as
@@
 Since for then we obtain
-Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [8] proved the following theorem.
+Let stand for the gradient operator on a Riemannian manifold. Cheng-Yau [8] proved the following theorem.
 Lemma 4.4.
@@
 Proof.
-By
-we obtain an estimate for upper bounds of as follows.
+Applying
+we obtain an upper bound estimate of as follows.
 Corollary 4.6.
@@
 Lemma 4.7 (Borel Lemma).
-Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
+Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that
 holds for all outside
@@
 In the following, we establish a logarithmic derivative lemma. Let be a meromorphic function on The norm of the gradient of is defined by
-in local holomorphic coordinates here is the inverse of Define the Nevanlinna’s characteristic function of by
+in a local holomorphic coordinate where is the inverse of Define the Nevanlinna’s characteristic function of by
 where
@@
 On one puts a singular metric
-which leads to
+with
 Lemma 4.9.
@@
 Note that
-Whence, it concludes from Fubini’s theorem that
+It follows from Fubini’s theorem that
+By this with the First Main Theorem, we conclude that
 Lemma 4.10.
@@
 Proof.
-Define a non-negative function on via
+Define a non-negative function on by
 where
 is a singular volume form on By we obtain
-in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1)
-On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and
+in the sense of currents, in which is the zero divisor of and is the ramification divisor of Using Dynkin formula, we have
+On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on and on as well as
 for On each write
@@
 Again, set Then on we have
-For any we can take local holomorphic coordinates near and a local holomorphic coordinate near such that
+For any we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
 It deduces that
@@
 By this with Dynkin formula and Theorem 4.11, for any there exists a subset of finite Lebesgue measure such that
-holds for all outside where one used the fact that Combining (4.2.4) and (19), we have the theorem proved. ∎
+holds for all outside where one used the fact that Combining (4.2.4) with (20), we have the theorem proved. ∎
 4.3. Second Main Theorem for Non-positive Sectional Curvature
@@
 A. Setups for
-Due to the curvature conditions, we have a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
+By the curvature conditions, there exists a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
 and
@@
 Here, is the standard Euclidean area of For define
-Since as and then we deduce that is a precompact domain containing in which exhausts So, the boundary of can be formulated as
+Since as and we deduce that is a precompact domain containing in which exhausts Thus, the boundary of can be formulated as
 Set
@@
 we can extend continuously to defined by (4).
-Notice that is parabolic when i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case.
+It is noted that is parabolic if i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case.
 However, we note (22) for it is thus natural to define
-for Let it then yields from (22) that
+for Let then it yields from (23) that
 To conclude, has the unified form
@@
 for
-Let be a meromorphic function. Similarly, one can defined the Nevanlinna’s functions and of as well as on as in Section 4.2.1 and Section 4.2.4.
 4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
@@
 Proof.
-When is the geodesic ball centered at with radius The conclusion holds immediately due to (23). In the following, we consider the other case. Since on then for
+When is the geodesic ball centered at with radius The conclusion holds immediately due to (4.3.1). In the following, we consider the other case. Write
+Since on we deduce that
+Hence, for
 This completes the proof. ∎
@@
 Proof.
-When is a geodesic ball centered at with radius Hence, it yields from (4) and (23) that
+When is a geodesic ball centered at with radius Hence, it yields from (4) and (4.3.1) that
 This implies the conclusion holds. For by Lemma 4.4 with (3), (4) and (21) (see Remark 5 in [34] also)
@@
 Substituting (28) into (4.3.2), we conclude that
-for outside , where ∎
+for outside with ∎
 Lemma 4.18.
@@
 This implies that the theorem holds. We consider the other case that Since as we deduce that
+and
 Thus, we conclude that
@@
 5. Nevanlinna Theory of Algebroid Functions
-Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
+Let be a complete non-compact Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
 5.1. First Main Theorem
-Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definitions in Section 4.1).
-In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
-Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus We can choose such that
+Let denote a precompact domain with smooth boundary in Fix a point As defined in Section 4.1, we have the Green function for and the harmonic measure on
+In what follows, we introduce Nevanlinna’s functions of algbroid functions. We are given a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
+such that defines a meromorphic mapping with indeterminacy locus As we note that are the distinct single-valued meromorphic components of on outside Assume that
 The characteristic function of for is defined by
@@
 Proof.
-With the aid of Vieta’s theorem, we obtain Dynkin formula yields that
-By definition, the left hand side equals
+According to the Vieta’s theorem, we obtain Combining Dynkin formula, it yields that
+By the definition, the left hand side equals
 On the other hand, Theorem 2.21 and Theorem 2.22 imply that the right hand side equals
+Then, we arrive at
+Setting to get
+where
+In particular, and Similarly, we obtain
+which leads to
+Since
 we obtain
-Set then
-with
-In particular, and Similarly, we have
-which means that
-Since
-then we obtain
 Put together the above, we conclude that
-5.2. Estimates of Branching Divisors
-Let be a -valued algebroid function on We use the same notations such as given in Section 5.1.
-Our aim is to give an estimate of the branching divisor of Denote by the -valued, -valued components of at any respectively.
-Then, we have namely, is a -order branching point of for and is hence a -order branching point of The branching divisor of associated to is defined by a divisor on satisfying that
+5.2. Estimate of Branching Divisor
+Let be a -valued algebroid function on We use the same notations as defined in Section 5.1.
+Our purpose is to obtain an estimate of the branching divisor of For each we use to denote the all -valued, -valued components of at respectively.
+As seen before, we have Namely, has branching order at with and has branching order at The branching divisor of associated to is defined to be the divisor on satisfying that
 The counting function of branching divisor of for is defined by
 where
-Review that the discriminant of is given by
+Recall that the discriminant of is given by
 Lemma 5.2.
 We have
-in the sense of currents.
 Proof.
-Take any Let be the -valued, -valued components of at respectively, where Since the set of singular branching points of is an analytic set of complex dimension not greater than then one just needs to consider that is a non-singular branching point of By Theorem 2.14, we can write for
-in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of So, one can derive from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.22, we have
-This implies the lemma holds. ∎
+For any we use to denote the all distinct function elements of at with leaves respectively.
+It is noted that From the assumption, we deduce that Since the set of all singular branching points of has complex codimension not smaller than it is sufficient to treat the case where is a non-singular branching point of By Theorem 2.14, we may write for
+in a local holomorphic coordinate near such that has a local holomorphic coordinate If then is not a pole of Through a simple observation of we deduce from (29) immediately that there exist terms that contain the factor i.e., terms that contain the factor in For the worst case that for all we deduce that the factor appears a total of at least times in Hence,
+If then is a pole of For the worst case where for all it yields from Theorem 2.21 that
+Combining the above two cases, we prove the lemma. ∎
 Set
-Theorem 5.3 (Estimates of branching divisors).
+Theorem 5.3 (Estimate of branching divisor).
 We have
 5.3. Second Main Theorem
-Let be the -leaf complex manifold of a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.21.
+Let be the -leaf complex manifold by a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.20.
 Let be the natural projection.
-Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula (see Lemma 4.1) still holds under the pullback metric
+Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula still holds under the pullback metric
 Set Picking a sequence of neighborhoods of such that
-Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then we may assume that (otherwise, a subsequence can be chosen so that as ).
+Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then one may assume that (otherwise, a subsequence can be chosen so that as ).
 Let be the Laplace-Beltrami operators of on respectively.
@@
 It is not difficult to show that as and (weakly) uniformally on as due to
-for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively.
-To unify the form, one uses to stand for the volume elements of for respectively.
-Apply Dynkin formula, it gives
+as for a Borel subset of Similarly, we have (weakly) uniformally on as in which stand for the volume measures of respectively. Set and Using Dynkin formula, we are led to
 for a -class function with on outside a polar set of singularities at most. Hence, it yields by letting that
-This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
+This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex codimension
 5.3.1. with non-negative Ricci curvature
@@
 Assume that We denote by the connected components of centered at respectively, with for Again, define
-Then, it is trivial to see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and also is the harmonic measure on with respect to Set
+Then, it is easy to infer that is the Green function of for with a pole at satisfying Dirichlet boundary condition, as well as that is the harmonic measure on with respect to Set
 where
@@
 Proof.
-Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (7), say where with and is a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
-On the other hand, we have and
-Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
-holds for all outside Hence, we have the theorem proved. ∎
-When we have By it is trivial to deduce Hence, we obtain:
+Note that is a -sheeted analytic covering, and two distinct connected components () may cross only at those points in the branching set of If is a -order branching point of then there are multi-leaf function elements with subordinate to equation (7).
+As we know that Since a point in may be a branching point of it yields from the definition of that
+On the other hand, we have as well as
+Combining the above with (5.3.1), it follows that for any there exists a subset of finite Lebesgue measure such that
+holds for all outside By Theorem 5.3, we prove the theorem. ∎
+When we have By it is trivial to deduce that Hence, we obtain:
 Corollary 5.5.
@@
 Assume that has non-positive sectional curvature satisfying (3).
-Note that the curvatures of also satisfy (3) under the pullback metric In the same way, we obtain notations and by setting Apply the similar arguments as in the proof of Theorem 5.4, we can show that
+Note that the curvature of also satisfies (3) under the pullback metric In the similar way, we obtain notations and by setting Applying the same arguments as in the proof of Theorem 5.4, we can show that
 Theorem 5.6 (Second Main Theorem).
@@
 holds for all outside
-If then and Hence, we conclude that
+When we deduce that and Thus, we conclude that
 Corollary 5.7.
@@
 Let be the unit ball centered at the origin 0 in Consider equipped with the Poincaré metric
-in holomorphic global coordinates where
-It is known that is a non-compact complete Kähler manifold of constant sectional curvature under this metric. From (29), the associated Laplace-Beltrami operator is
+in a global holomorphic coordinate where
+It is known that is a complete non-compact Kähler manifold of constant sectional curvature under this metric. From (31), the associated Laplace-Beltrami operator is
 Take as a reference point. The associated Poincaré distance function of from 0 on can be easily written as
@@
 is the standard Euclidean metric on It yields from (31) that
-where
-Let be the Riemannian area element of Using the second formula in (31) and volume form formula, it gives immediately
+with Let be the Riemannian area element of By the second formula in (33) and volume form formula, it gives immediately
 where is the standard Euclidean area element of
-With these preparations, we shall give an estimate for Ricci term Let be the scalar curvature of which is a negative constant. It follows from (32) that
-Substituting (30) into the expression of and calculating the integral, it is not hard to deduce that
+With these preparations, we shall give an upper estimate of Let be the scalar curvature of which is a negative constant. It yields from (34) that
+Substituting (32) into the expression of and computing the integral, it is not hard to deduce that
 We proceed to estimate with By
@@
 6. Unicity Problems of Algebroid Functions
-Assume that is a non-compact complete Kähler manifold.
+Assume that is a complete non-compact Kähler manifold.
 6.1. Propagation of Algebraic Dependence
@@
 Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable.
-Given algebroid functions Put
+Given algebroid functions Set
 Definition 6.1.
 Let be an analytic subset of We say that nonconstant algebroid functions on are algebraically dependent on if there exists a proper indecomposable algebraic subset of such that If so, are said to be -related on
-Let be an effective divisor on and fix any positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
-It is clear that Set
-Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ).
-We introduce the notation as follows.
-When is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), denote by
+We let be an effective divisor on and fix a positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
+Note that Set
+Let be hypersurfaces of satisfying that if or if for Let be distinct values in and let be positive integers (allowed to be ).
+We introduce the notation as follows:
+If is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), then denote by
 the set of all such that
-When has non-positive sectional curvature satisfying (3), denote by
-the set of all such that
-and (5) is satisfied.
-Let be the hyperplane line bundle over It gives a holomorphic line bundle over
-where is the natural projection on the -th factor with Let for For , define
+Let be the point line bundle over It gives a holomorphic line bundle over :
+where is the natural projection on the -th factor with Let with For , define
 with
-where is a fixed positive -form.
+where is a positive -form.
 Lemma 6.2.
@@
 Thus, we have
-Apply Dynkin formula, we deduce that
+Applying Dynkin formula, we deduce that
 This completes the proof. ∎
-Let be the set of all indecomposable hypersurfaces in such that for some Put
+We denote by the set of all indecomposable hypersurfaces of such that for some Set
 Lemma 6.3.
@@
 Proof.
-By we obtain
-in the sense of currents, where and are understood as divisors. Thus, we obtain
+Since we obtain
+in the sense of currents, where and are understood as divisors. Thus, we are led to
 By this with Theorem 5.4, we obtain
@@
 Set
-and
 where
@@
 Proof.
-Put
+Set
 Fix any -valued algebroid function We say that is generic with respect to if for at least one Assume that is generic with respect to Let be the set of all such that on Set
@@
 Let be a -valued algebroid function in It suffices to show that on Set
-Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the hyperplane line bundle over If on then there is a holomorphic section of over such that and Taking then we obtain and On the other hand, implies that
+Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the point line bundle over If on then there is a holomorphic section of over such that and Taking then one can have and On the other hand, implies that
 where With the help of Theorem 6.5, we obtain which is a contradiction. This proves the theorem. ∎
@@
 Assume on the contrary that The condition implies that
-in the sense of currents, where and are understood as divisors. So, we are led to
-It yields from Lemma 6.3 that
-Similarly,
-Combining (6.2) with (6.2), it yields from Theorem 5.4 that
+in the sense of currents, where and are understood as divisors. Thus, we obtain
+In further, it follows from Lemma 6.3 that
+Similarly, we have
+Combining the above and using Theorem 5.4, we are led to
 It is therefore

### v11

@@
 - 2 Algebroid Functions on Hermitian Manifolds
-- 3 Uniformization of Algebroid Functions
+- 3 Uniformization of Algebroid Functions without Construction
 - 4 Nevanlinna Theory of Meromorphic Functions
@@
 Darboux deemed that it is an important class of functions.
-To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a generic -sheeted analytic covering of A further fact states that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
+To see that, for example, it is known that a -valued algebroid function defines a -leaf Riemann surface which is a -sheeted ramified analytic covering of A further fact states that can lift to a single-valued meromorphic function on via a natural projection such that which is called the uniformization of
 The most concerned problem on algebroid functions seems to be the value distribution.
@@
 We generalize the notion of algebroid functions from to (see Definition 2.15).
-Let be a -valued algebroid function on Given a -leaf function element of When is not a singular branching point of we give a local Pusieux expansion of at i.e., in a local holomorphic coordinate we show that can be expanded locally into the following convergent Pusieux series (see Theorem 2.14)
-at with where are meromorphic functions. By this Pusieux series, we construct a -leaf complex manifold and prove that
+Let be a -valued algebroid function on Given a -leaf function element of When is not a singular branch point of we give a local Pusieux expansion of at i.e., in a local holomorphic coordinate we prove that can be expanded locally into the following convergent Pusieux series (see Theorem 2.14)
+at with where are meromorphic functions. By this Pusieux series, we construct a -leaf complex manifold and show that
 Theorem A.
@@
 Note that as due to
-The condition (2) is relaxed. For example, it holds when with and holds when with or and
+The condition (2) is relaxed. For example, it is satisfied when with or when with or and
 Corollary B3 (Picard Type Theorem).
@@
 (b2) has non-positive sectional curvature
-Assume that is non-positively curved satisfying
+Assume that has curvature satisfying
 for constants with that either or where are the sectional curvature and Ricci curvature of respectively. Set
@@
 Stoll [36], etc.
-Assume that is either non-parabolic with non-negative Ricci curvature, or has non-positive sectional curvature satisfying (3).
+Assume that is either non-parabolic with non-negative Ricci curvature, or of curvature satisfying (3).
 Let be an effective divisor on where are irreducible hypersurfaces of Let be a positive integer (allowed to be ). Set
@@
 be the set of all such that
-and (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has non-positive sectional curvature satisfying (3).
+and that either (2) is satisfied if is non-parabolic with non-negative Ricci curvature, or (5) is satisfied if has curvature satisfying (3).
 Let be the point line bundle over We denote by the set of all indecomposable hypersurfaces of such that for some in which and is the natural projection on the -th factor. Set
@@
 (d1) has non-negative Ricci curvature
-Assume that is non-parabolic satisfying (2). Let be a positive integer. Consider the following condition
+Assume that is non-parabolic with volume growth satisfying (2). Let be a positive integer. Consider the following condition
 with
@@
 (d2) has non-positive sectional curvature
-Assume that is non-positively curved satisfying (3).
+Assume that has curvature satisfying (3).
 Theorem D3.
@@
 where are holomorphic functions locally defined on such that they well define a meromorphic mapping
-Let be the indeterminacy locus of which is either an analytic set of complex codimension not smaller than for or an empty set for in Then, is holomorphic on
+Let be the indeterminacy locus of which is either an analytic set of complex codimension not smaller than for or an empty set for in Note that is holomorphic on
 Remark 2.1.
-The local definition for means that a collection of holomorphic functions i.e., is a holomorphic function for all satisfying that
+The local definition for means that i.e., a collection of holomorphic functions for all such that
 where is an open covering of Whence, we see that is well defined on and it hence defines a meromorphic function on with
@@
 The symmetry implies that can also continue analytically to along However, we should mention that won’t be, in general, continued analytically back to along respectively.
-In what follows, we would consider the analytic continuation of a function element along a closed Jordan curve surrounding any multiple point Take a smooth closed Jordan curve around such that which is close to and doesn’t surround or cross any connected components of that do not pass through Without loss of generality, we look at the analytic continuation of along starting from When returns to along it will turn to certain function If then one continues to extend analytically along in the same direction of rotation.
-Repeating this action, we shall see that cycles on and on periodically.
-Let be the smallest period such that can return to itself along all such curves Then, we get We call a branching point with order of for Repeatedly, it can be seen that is divided into finitely many groups.
-Let denote the number of groups.
-If the -th group has members, then
+In what follows, we would consider the analytic continuation of a function element along a closed Jordan curve surrounding any multiple point Take a smooth closed Jordan curve around such that which is close to and doesn’t surround or cross any connected components of that do not pass through Without loss of generality, we look at the analytic continuation of along starting from When returns to along it will turn to certain function by which we mean that for a local coordinate chart ( is small enough), jumps to some single-valued component when travels around once along started at due to the change of -dimensional argument of If then one continues to extend analytically along in the same direction of rotation.
+Repeat this action, we shall see finally that cycles on and on periodically.
+Let be the smallest period such that can return to itself along all such curves Then, we can obtain We call a branch point with order of if Repeatedly, it can be seen that is divided into finitely many groups.
+Let be the number of groups.
+If the -th group has members for then
 We just analyze the first group: Taking a suitable there exists a biholomorphic mapping such that
-where is the polydisc centered at 0 with polyradius in Setting where is the unit disc centered at 0 in the -plane. For any we have Set and write
-We show that is a single-valued analytic function on Notice that is the branching order of we see that is holomorphic on because is so close to 0 that there exist no any branching points of in We only need to examine that is a removed singularity of To do so, it suffices to prove that is bounded near 0 due to Riemann’s analytic continuation.
+where is the polydisc centered at 0 with polyradius in Setting where is the unit disc centered at 0 in the -plane. For any we have
+Set and write
+We show that is a single-valued analytic function on Notice that is the branch order of we deduce that is holomorphic on because is so close to 0 that there exist no any branch points of in We only need to examine that is a removed singularity of To do so, it suffices to prove that is bounded near 0 due to Riemann’s analytic continuation.
 In fact, it follows from (7) and (9) that
 on If we obtain Thus, on i.e., is bounded on It means that is holomorphic on and it has the Taylor expansion
-at Let be the set of all branching points of in Say that a branching point of is non-singular, if does not lie in a crossing of components of and singular otherwise.
-Assume that is not a singular branching point of Then, is smooth if is small enough.
+at Let be the set of all branch points of in Saying that a branch point of is non-singular, if does not lie in any crossings of components of and singular otherwise.
+Assume that is not a singular branch point of Then, is smooth when is small enough.
 Taking a suitable if necessary.
-Since is bounded on and defines the branching order of at which is independent of the choice of one can deduce from and Riemann’s analytic continuation that there exists a biholomorphic mapping with
+Since is bounded on and defines the branch order of at that is independent of the choice of we can conclude from and Riemann’s analytic continuation that there exists a biholomorphic mapping with
 such that can be expanded locally into the following Pusieux series
@@
 According to the above arguments, we show that is a -valued analytic function defined in a neighborhood of We call a -leaf function element.
-If one cuts along a real hypersurface passing through then can separate into single-valued analytic components, and all these components can turn to each other when they continue analytically along a closed curve around in that is sufficiently close to By the local expansion of we deduce that is a -multiple, -valued branching point with order of
-Assume that is not a singular branching point of Set and define
+If one cuts along a real hypersurface passing through then can separate into single-valued analytic components, and all these components can turn to each other when they continue analytically along a closed curve around in that is sufficiently close to By the local expansion of we note that is a -multiple, -valued branch point with order of
+Assume that is not a singular branch point of We set and define
 By the previous arguments, one can also expand locally into a Pusieux series of the form
@@
 Hence, the case turns to the case
-To conclude, for each there are distinct multi-leaf function elements subordinate to equation (7), denoted by where is a -valued meromorphic function with and Hence, has branching order at with When we say that is a branching point with order of Moreover, if is not a singular branching point of then we deduce that can expand locally into the convergent Pusieux series at That is, we have shown that
+To conclude, for each there are distinct multi-leaf function elements subordinate to equation (7), denoted by where is a -valued meromorphic function with and Hence, has branch order at with If then we say that is a branch point with order of Furthermore, if is not a singular branch point of then we deduce that can expand locally into the convergent Pusieux series at That is, we have shown that
 Theorem 2.14.
-Let be a -leaf function element subordinate to equation If is not a singular branching point of then can be expanded locally into the following convergent Pusieux series
-at in a local holomorphic coordinate such that and for all where is the branching set of and are meromorphic functions.
+Let be a -leaf function element subordinate to equation If is not a singular branch point of then can be expanded locally into the following convergent Pusieux series
+at in a local holomorphic coordinate such that and for all where is the branch set of and are meromorphic functions.
 Note that in Theorem 2.14 is a -multiple root of that is, can separate into distinct single-valued components with value at If then separates into distinct single-valued components which take at
-According to the analytic continuation and uniqueness theorem, the above arguments show that equation (7) uniquely defines a -valued meromorphic function on outside the branching points of Let denote the set of all branching points of which is an analytic set in called the branching set of We generalize the notion of algebroid functions as follows.
+According to the analytic continuation and uniqueness theorem, the above arguments show that equation (7) uniquely defines a -valued meromorphic function on outside the branch points of Let stand for the set of all branch points of which is an analytic set in called the branch set of We extend the notion of algebroid functions from to as follows.
 Definition 2.15.
@@
 where are holomorphic functions locally defined on such that they define a meromorphic mapping
-In a word, denote by the distinct single-valued meromorphic components of and given a -leaf function element subordinate to equation (7), with distinct single-valued components Then, there is a permutation of such that on
-In the end, we provide an example to illustrate how to calculate the order of branching points of algebroid functions.
-Example. Calculating the order of the -valued algebroid function
-on at branching points
+In a word, denote by the distinct single-valued components of and given a -leaf function element subordinate to equation (7), with distinct single-valued components Then, there exists a permutation of with for on
+In the end, we provide an example to illustrate how to calculate the order of branch points of algebroid functions.
+Example. Calculating the branch order of the algebroid function
+on at branch points
 Since have single-valued components, respectively, we infer that has single-valued components; Since have single-valued components, respectively, we infer that has single-valued components.
 Thus, is a -valued algebroid function.
-Let stand for any closed Jordan curve in All the branching sets of are as follows
+Let stand for any closed Jordan curve in All the branch sets of are as follows
 It is clear that
 Order of
-Take such that is surrounded by Note that are the only two branching sets of that pass through Therefore, is not allowed to surround or pass through any other branching sets of except for in order to determine the order of For all such we for
-Thus, the order of is determined by the order of branching point of and the order of branching point of It is
+Take such that is surrounded by Note that are the only two branch sets of that pass through So, is not allowed to surround or pass through any other branch sets of except for in order to determine the order of For all such we for
+Hence, the order of is determined by the orders of branch point of and the order of branch point of It is
 Order of
-Take such that is surrounded by Note that are the only two branching sets of that pass through Similarly, cannot surround or pass through any other branching sets of except for It implies that for
-Thus, the order of is determined by the order of branching points of It is
+Take such that is surrounded by Note that are the only two branch sets of that pass through Similarly, cannot surround or pass through any other branch sets of except for It implies that for
+Hence, the order of is determined by the orders of branch points of It is
 Order of
-Take such that is surrounded by Since are the only three branching sets of that pass through we have for
-Thus, the order of is determined by the order of branching points of respectively. It is
+Take such that is surrounded by Since are the only three branch sets of that pass through we have for
+Hence, the order of is determined by the orders of branch points of respectively. It is
 Order of
-Take such that is surrounded by Since are the only three branching sets of that passes through we have for
-Thus, the order of is determined by the order of branching points of It is
+Take such that is surrounded by Since are the only three branch sets of that passes through we have for
+Hence, the order of is determined by the orders of branch points of It is
 2.2. Multi-leaf Complex Manifolds by Algebroid Functions
@@
 We define an equivalent relation as follows: if and only if
-for some two single-valued components of respectively, where is the branching set of with Set
+for some two single-valued components of respectively, where is the branch set of with Set
 Definition 2.18.
-A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one may regard that
+A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branch set of If so, one may regard that
 Let A neighborhood of is a set which contains the following elements:
@@
 Next, we will equip with a complex structure such that is a complex manifold.
-Taking a -leaf function element Assume that is not a singular branching point of and is chosen suitably.
-Thanks to Theorem 2.14, there exists a biholomorphic mapping such that can be expanded into the Pusieux series
+Taking a -leaf function element Assume that is not a singular branch point of Let be chosen suitably.
+By means of Theorem 2.14, there exists a biholomorphic mapping such that can be expanded into the Pusieux series
 on with and for all
@@
 where is an integer such that
-Let be the set of all function elements satisfying that is a singular branching point of Moreover, denotes the set of all the corresponding singular branching points.
-It is clear that is an analytic set of complex codimension not smaller than Set
+Let stand for the set of all function elements such that is a singular branch point of Moreover, we denote by the set of all the corresponding singular branch points of which is obvious an analytic set of complex codimension not smaller than Set
 We now show that is a complex atlas of Assume that It is sufficient to verify that the transition mapping
@@
 where is the -th single-valued component of Evidently, is also a holomorphic mapping. Hence, is a non-compact -leaf complex manifold of complex dimension
-Treat the singular part Let be the natural projection by mapping any function element to Note that is a generic -sheeted analytic covering such that
+Finally, we shall treat the singular branch set Let be the natural projection which maps a function element to Note that is a -sheeted ramified analytic covering such that
 Since is an analytic set of complex codimension not smaller than the complex structure of can be viewed as a natural extension of the complex structure of This implies that one can naturally extend the complex structure of to via Hence, we conclude that
 Theorem 2.20.
-is a non-compact -leaf complex manifold of complex dimension We call the complex manifold by
+is a -sheeted ramified analytic covering of We call the -leaf complex manifold by
+Corollary 2.21 (Uniformization).
+An algebroid function on can lift to a meromorphic function on via the natural projection by mapping to such that
 2.3. Zero and Pole Divisors of Algebroid Functions
-Let be a -valued algebroid function on defined by equation Let be single-valued meromorphic components of As a matter of convenience, we write
+Let be a -valued algebroid function on defined by equation Let be single-valued components of
 Theorem 2.22.
@@
 Let be a zero of Since in the sense of divisors, we may assume that It follows that
-where denotes the number of -valued points of a function at Without loss of generality, suppose that are all single-valued components such that which take at and suppose that are all single-valued components such that which take at Then, we have By Vieta’s theorem, it yields that
-Combining this with (11), we see that reaches the maximal order of a pole of Since we obtain
-which means that ∎
+where denotes the number of -valued points of a function at Without loss of generality, we assume that are all components that take at According to Vieta’s theorem, we obtain
+It implies that since reaches the maximal pole order at It yields from (11) that
+Thus, we conclude that ∎
 Similarly, we can show by setting that
@@
 We have
-3. Uniformization of Algebroid Functions
-This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold.
-As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold.
-The main idea originates from W.
-Stoll [35].
+3. Uniformization of Algebroid Functions without Construction
+In this section, we provide another approach to realize the uniformization of algebroid functions without constructing finite ramified analytic coverings. Our main idea originates from W. Stoll [35].
 3.1. Algebroid Reduction
-Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, by which one means a proper, surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex codimension It is clear that
-is a proper, surjective, local biholomorphic mapping of sheet number and thus a -sheeted analytic covering. By R. Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
+Let be a non-compact complex manifold of complex dimension Let be a -sheeted ramified analytic covering, which means a proper, surjective holomorphic mapping with a ramification set i.e.,
+is a proper, surjective, local biholomorphic mapping of sheet number and thus a -sheeted analytic covering.
+Here, is the set of those points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex codimension By R.
+Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
 is an analytic set satisfying that and Moreover, and for Let The mapping degree of at is defined in such manner: if is a neighborhood of such that then
@@
 is called the symmetric tensor product of and stands for the -order symmetric group. Note that is a subspace of -fold tensor product space with Again, put
-where for any Naturally, is embedded into by the identity mapping Thus, we obtain
+where for any Naturally, is embedded into by the mapping Thus, we obtain
 Definition 3.1.
@@
 3.2. Uniformization
-In what follows, we consider the case that For a convenience, we write and in short. Take a basis of Then for any it has a unique representation
-for some numbers It gives a natural isomorphism defined by
-which yields that If then (13) defines an algebraic equation
+In what follows, we consider the case when For convenience, we write and for short. Take a basis of For any there is a unique representation
+for certain It induces a natural isomorphism defined by
+which follows that If then (13) defines an algebraic equation
 Let be the maximal integer in such that Then, equation (14) reduces to Let be the roots of the equation. Applying Vieta’s theorem, it is immediate that
@@
 Theorem 3.7 (Uniformization).
-Let be a -valued algebroid function on Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
+Let be an algebroid function. Then, there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a meromorphic function such that
 Proof.
@@
 Set with a basis By (15), we can write
-in the sense of isomorphisms. Further, the Vieta’s theorem leads to
+in the sense of isomorphisms. In further, the Vieta’s theorem leads to
 in the sense of isomorphisms.
@@
 with
-Let be -valued points of Then
+Let be the -valued points of Then
 Combining with then we prove the theorem. ∎
-In conclusion, given a -leaf complex manifold by a -valued algebroid function on there exists the natural projection of generic sheet number given by which is a proper surjective holomorphic mapping.
-That is, we have shown that
-Corollary 3.8.
-Let be the -leaf complex manifold by a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
 4. Nevanlinna Theory of Meromorphic Functions
@@
 where is the Riemannian volume element of and is the inward normal derivative on
-Next, we introduce Nevanlinna’s functions. Let be a meromorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
-The characteristic function of for is defined by
-Let The proximity function of for with respect to is defined by
+Next, we introduce Nevanlinna’s functions. Let be a meromorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and with
+The characteristic function of on with respect to is defined by
+Let The proximity function of on with respect to is defined by
 where denotes the spherical distance on defined by
 with
-The counting function of for with respect to is defined by
-Lemma 4.1 (Dynkin Formula, [1, 11, 12]).
+The counting function of on with respect to is defined by
+We need the following Dynkin formula (see, e.g., [1, 11, 12]).
+Lemma 4.1 (Dynkin Formula).
 Let be a -class function on outside a polar set of singularities at most. Assume that Then
@@
 Proof.
-Applying Dynkin formula to we obtain
+Applying Dynkin formula to we get
 Note that the first term on the left hand side equals and the right hand side equals
+This proves the theorem. ∎
 4.2. Second Main Theorem for Non-negative Ricci Curvature
-Assume that is non-parabolic with non-negative Ricci Curvature.
-Fix a reference point and denote by the geodesic ball centered at with radius in Let denote the Riemannian volume of The non-parabolicity of means that
-Thus, we have the unique minimal positive global Green function of for which can be written as
-where is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
+Assume that is non-parabolic. Fix a reference point and denote by the geodesic ball centered at with radius in Let denote the Riemannian volume of The non-parabolicity of means that
+Thus, we have the unique minimal positive global Green function of for We can write as
+in which is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
 holds for all
@@
 For define
-Since and then we conclude that is a precompact domain containing satisfying that and In fact, the family exhausts i.e., for any sequence with we have
+Since
+we see that is a precompact domain containing with and Moreover, the family can exhaust i.e., for any sequence with we have
 So, the boundary of can be formulated as
@@
 Let be a meromorphic function. Take in Section 4.1, we have the Nevanlinna’s functions of :
-4.2.2. Gradient estimates of Green functions
+4.2.2. Gradient estimate of Green functions
 Lemma 4.3.
@@
 Proof.
-According to the definition of Green function for it is immediate that for
-Since for then we obtain
+Invoking the definition of Green function for it is immediate to deduce that for
+Since for we obtain
 Let stand for the gradient operator on a Riemannian manifold. Cheng-Yau [8] proved the following theorem.
@@
 4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
-Lemma 4.7 (Borel Lemma).
+Lemma 4.7 (Borel’s Growth Lemma).
 Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that
@@
 In further, we have
-Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
+Applying Borel’s growth lemma to the left hand side of this equality twice: one is to and another is to we deduce that for any there exists a subset of finite Lebesgue measure such that
 holds for all outside On the other hand, Corollary 4.6 implies that there exists a constant such that
@@
 For any we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
-It deduces that
+It yields that
 Put Then, we obtain
-on Therefore,
+on It is therefore
 on Since is bounded on then it yields that
 By this with Dynkin formula and Theorem 4.11, for any there exists a subset of finite Lebesgue measure such that
-holds for all outside where one used the fact that Combining (4.2.4) with (20), we have the theorem proved. ∎
+holds for all outside in which we used the fact that Combining (4.2.4) with (20), we have the theorem proved. ∎
 4.3. Second Main Theorem for Non-positive Sectional Curvature
-Assume that has non-positive sectional curvature satisfying (3).
-From the definition of curvatures, it is not difficult to see that In particular, we have if has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering As long as is equipped with the pull-back metric induced from via then has the same curvatures as ones of In fact, is a Cartan-Hadamard manifold.
-In the following, we denote by the Riemannian distance function of from on
+Assume that has curvature satisfying (3).
+According to the definition of curvature, it is not difficult to see that In particular, we have when has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering If is equipped with pullback metric induced from via then preserves the same curvature as Note that is a Cartan-Hadamard manifold.
+In the following, we use to denote the Riemannian distance function of from on
 4.3.1. Construction of and Nevanlinna’s functions
@@
 Here, is the standard Euclidean area of For define
-Since as and we deduce that is a precompact domain containing in which exhausts Thus, the boundary of can be formulated as
+Since as and one infers that is a precompact domain containing in which exhausts Thus, the boundary of can be formulated as
 Set
@@
 There exists a constant such that
-holds for
+holds for all
 Proof.
@@
 for some large constant which depends only on the dimension By (21) again
-which gives for So, we conclude that
+which implies that for all Thus, we conclude that
 where ∎
@@
 There exists a constant such that
-holds for where is the Riemannian area element of
+holds for all where is the Riemannian area element of
 Set
@@
 Let be a locally integrable function on Assume that is locally bounded at Then for any there exists a set of finite Lebesgue measure such that
-holds for outside where is defined by
+holds for all outside where is defined by
 Proof.
 Lemma 4.14 yields that
-Differentiating to obtain
-In further
-Applying Borel’s lemma to the left hand side on the above equality twice, then we see that there exists a set of finite Lebesgue measure such that
-holds for outside Meanwhile, Corollary 4.16 means that there exists a constant such that
+Differentiating to get
+In further, we have
+Applying Borel’s growth lemma to the left hand side on this equality twice, we can deduce that there exists a set of finite Lebesgue measure such that
+holds for all outside Meanwhile, Corollary 4.16 means that there exists a constant such that
 Substituting (28) into (4.3.2), we conclude that
-for outside with ∎
+for all outside with ∎
 Lemma 4.18.
 Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
-holds for outside where is defined by
+holds for all outside where is defined by
 Proof.
@@
 Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
-holds for outside where is defined by
+holds for all outside where is defined by
 Proof.
@@
 Let denote a precompact domain with smooth boundary in Fix a point As defined in Section 4.1, we have the Green function for and the harmonic measure on
-In what follows, we introduce Nevanlinna’s functions of algbroid functions. We are given a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
-such that defines a meromorphic mapping with indeterminacy locus As we note that are the distinct single-valued meromorphic components of on outside Assume that
-The characteristic function of for is defined by
+In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branch set on which is defined by an irreducible algebraic equation
+such that defines a meromorphic mapping with indeterminacy locus In the following, we assume that
+The characteristic function of on is defined by
 where
-are called the proximity function, counting function of for respectively. For define
+are called the proximity function, counting function of on respectively. For define
 where
@@
 Proof.
-According to the Vieta’s theorem, we obtain Combining Dynkin formula, it yields that
+According to Vieta’s theorem, we have
+Since is polar, it yields from Dynkin formula that
 By the definition, the left hand side equals
@@
 where
-In particular, and Similarly, we obtain
-which leads to
+In particular, we have
+Similarly, we obtain
+which yields that
 Since
-we obtain
+we deduce that
 Put together the above, we conclude that
-5.2. Estimate of Branching Divisor
+5.2. Estimate of Branch Divisors
 Let be a -valued algebroid function on We use the same notations as defined in Section 5.1.
-Our purpose is to obtain an estimate of the branching divisor of For each we use to denote the all -valued, -valued components of at respectively.
-As seen before, we have Namely, has branching order at with and has branching order at The branching divisor of associated to is defined to be the divisor on satisfying that
-The counting function of branching divisor of for is defined by
+Our purpose is to obtain an estimate of the branch divisor of For each we use to stand for the all -valued, -valued components of at respectively.
+Then, we have and has branch order at with and has branch order at The branch divisor of associated to is defined by a divisor on such that
+The counting function of branch divisor of for is defined by
 where
@@
 For any we use to denote the all distinct function elements of at with leaves respectively.
-It is noted that From the assumption, we deduce that Since the set of all singular branching points of has complex codimension not smaller than it is sufficient to treat the case where is a non-singular branching point of By Theorem 2.14, we may write for
-in a local holomorphic coordinate near such that has a local holomorphic coordinate If then is not a pole of Through a simple observation of we deduce from (29) immediately that there exist terms that contain the factor i.e., terms that contain the factor in For the worst case that for all we deduce that the factor appears a total of at least times in Hence,
+It is noted that From the assumption, we deduce that Since the set of all singular branch points of has complex codimension not smaller than it is sufficient to deal with the case where is a non-singular branch point of By Theorem 2.14, we may write for
+in a local holomorphic coordinate near such that has a local holomorphic coordinate If then is not a pole of Through a simple observation of we deduce from (29) immediately that there exist terms that contain the factor i.e., terms that contain the factor in For the worst case where for all we deduce that the factor appears a total of at least times in Hence, we get
 If then is a pole of For the worst case where for all it yields from Theorem 2.22 that
@@
 Set
-Theorem 5.3 (Estimate of branching divisor).
+Theorem 5.3 (Estimate of branch divisors).
 We have
 5.3. Second Main Theorem
-Let be the -leaf complex manifold by a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.20.
+Let be the -leaf complex manifold by a -valued algebroid function on which is of complex dimension by Theorem 2.20.
 Let be the natural projection.
-Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula still holds under the pullback metric
+We equip with the pullback metric induced from via Note that is a positive semi-definite Kähler metric on i.e., is a Kähler metric on and degenerate on Since is a polar set with Dynkin formula still works on for the following reasons.
 Set Picking a sequence of neighborhoods of such that
-Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then one may assume that (otherwise, a subsequence can be chosen so that as ).
-Let be the Laplace-Beltrami operators of on respectively.
-We have weakly as Take an arbitrary bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and again by the harmonic measures of on with respect to respectively.
-It is not difficult to show that as and (weakly) uniformally on as due to
-as for a Borel subset of Similarly, we have (weakly) uniformally on as in which stand for the volume measures of respectively. Set and Using Dynkin formula, we are led to
-for a -class function with on outside a polar set of singularities at most. Hence, it yields by letting that
-This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex codimension
-5.3.1. with non-negative Ricci curvature
+Let be a partition of the unity subordinate to the open covering of Then, and Fix a Kähler metric on with Kähler form Set Since is positive definite on we see that is also positive definite on Moreover, since we conclude that is a Kähler metric on with on Let denote the Riemannian measures of respectively.
+By and as we can deduce that weakly as Let stand for the Laplace-Beltrami operators of on respectively.
+According to and weakly as one has weakly as Taking a bounded domain with smooth boundary in Fixing with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, here is well-defined on Also, we denote by the harmonic measures of on with respect to respectively, here is well-defined on Since is a polar set, the previous arguments can show that in the sense of distributions and weakly as By Dynkin formula
+for a -class function with on outside a polar set of singularities at most. Since is polar, we conclude by letting that
+This means that Dynkin formula works on under the pullback metric By Corollary 2.21, one can lift to a meromorphic function such that In further, it is not difficult to examine that Theorem 4.2, Theorem 4.12 and Theorem 4.20 can apply to on each leaf of under the corresponding conditions.
+5.3.1. has non-negative Ricci curvature
 Assume that is non-parabolic.
-With the pullback metric is also non-parabolic, with non-negative Ricci curvature.
-Let be the Kähler form of Take where is a bounded domain defined in Section 4.2.1.
-Recall the Green function for and the harmonic measure on defined in Section 4.2.1.
+With the pullback metric one notes that is non-parabolic, with non-negative Ricci curvature outside Let be the Kähler form of Set where is the bounded domain defined in Section 4.2.1.
+We recall the definitions of Green function for and harmonic measure on in Section 4.2.1.
 Then, the characteristic function of is defined by
@@
 The simple counting function of with respect to is defined by
-Assume that We denote by the connected components of centered at respectively, with for Again, define
-Then, it is easy to infer that is the Green function of for with a pole at satisfying Dirichlet boundary condition, as well as that is the harmonic measure on with respect to Set
+Assume that Let be the connected components of centered at respectively, with that for each Let denote the Laplace-Beltrami operator of which is well-defined on Define
+In a sense, we can regard outside that is the Green function of for with a pole at satisfying Dirichlet boundary condition, and that is the harmonic measure on with respect to Set
 where
@@
 Proof.
-Note that is a -sheeted analytic covering, and two distinct connected components () may cross only at those points in the branching set of If is a -order branching point of then there are multi-leaf function elements with subordinate to equation (7).
-As we know that Since a point in may be a branching point of it yields from the definition of that
+Note that is a -sheeted analytic covering, and two distinct connected components of meet only at the points in Moreover, any point in is possibly a branch point of or equivalently, any point in is possibly a ramification point of Hence, it follows from that
 On the other hand, we have as well as
-Combining the above with (5.3.1), it follows that for any there exists a subset of finite Lebesgue measure such that
+Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
 holds for all outside By Theorem 5.3, we prove the theorem. ∎
@@
 holds for all outside
-5.3.2. with non-positive sectional curvature
-Assume that has non-positive sectional curvature satisfying (3).
-Note that the curvature of also satisfies (3) under the pullback metric In the similar way, we obtain notations and by setting Applying the same arguments as in the proof of Theorem 5.4, we can show that
+5.3.2. has non-positive sectional curvature
+Assume that has curvature satisfying (3).
+Note that the curvature of also satisfies (3) outside under the pullback metric Similarly, we obtain notations and by setting Employing the similar arguments as in the proof of Theorem 5.4, we can conclude from Theorem 4.20 that
 Theorem 5.6 (Second Main Theorem).
@@
 in a global holomorphic coordinate where
-It is known that is a complete non-compact Kähler manifold of constant sectional curvature under this metric. From (31), the associated Laplace-Beltrami operator is
+It is known that is a complete non-compact Kähler manifold of constant sectional curvature under this metric. Using (31), the associated Laplace-Beltrami operator is
 Take as a reference point. The associated Poincaré distance function of from 0 on can be easily written as
-Hence, we have the minimal positive global Green function of for with a pole at 0:
-where denotes the standard Euclidean area of the unit sphere in By definition (see Section 2.1), we have
+Hence, we obtain the minimal positive global Green function of for with a pole at 0:
+where stands for the standard Euclidean area of the unit sphere in By the definition (see Section 2.1), we have
 This gives
 which is just the geodesic ball centered at 0 with radius in Hence, the boundary of is the geodesic sphere centered at 0 with radius Whence, we obtain the Green function of for with a pole at 0 satisfying Dirichlet boundary condition:
-So, for with
+Thus, for with
 We write
@@
 is the standard Euclidean metric on It yields from (31) that
-with Let be the Riemannian area element of By the second formula in (33) and volume form formula, it gives immediately
+with Let be the Riemannian area element of By the second formula in (33), it gives immediately
 where is the standard Euclidean area element of
 With these preparations, we shall give an upper estimate of Let be the scalar curvature of which is a negative constant. It yields from (34) that
-Substituting (32) into the expression of and computing the integral, it is not hard to deduce that
-We proceed to estimate with By
+Substituting (32) into the expression of and computing the integral, it is not hard to deduce that We proceed to estimate with By
 we have Hence, we obtain:
@@
 We introduce the notation as follows:
-If is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), then denote by
+If is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), then we denote by
 the set of all such that
@@
 Thus, we have
-Applying Dynkin formula, we deduce that
-This completes the proof. ∎
+Applying Dynkin formula, we conclude that This completes the proof. ∎
 We denote by the set of all indecomposable hypersurfaces of such that for some Set
@@
 Proof.
-Let be a -valued algebroid function in It suffices to show that on Set
-Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the point line bundle over If on then there is a holomorphic section of over such that and Taking then one can have and On the other hand, implies that
+Let be a -valued algebroid function in It suffices to show that on Set Let denote the diagonal of Since on we obtain Let be the natural projection on the -th factor with We consider the holomorphic line bundle over where is the point line bundle over If on then there exists a holomorphic section of over such that and Take then one can have and On the other hand, implies that
 where With the help of Theorem 6.5, we obtain which is a contradiction. This proves the theorem. ∎
@@
 A1 has non-negative Ricci curvature
-Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
+Assume that is non-parabolic with volume growth satisfying (2).
 Theorem 6.7.
@@
 Assume for all If then
-In particular, the conclusions hold for nonconstant if
 Proof.
@@
 By Theorem 6.6, we have holds. Similarly, when for we deduce that
-which implies holds. ∎
+which implies that holds. ∎
 Corollary 6.8 (Five-Value Type Theorem).
@@
 B1 has non-positive sectional curvature
-Assume that has non-positive sectional curvature satisfying (3).
+Assume that has curvature satisfying (3).
 Theorem 6.9.
@@
 A2 has non-negative Ricci curvature
-Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
+Assume that is non-parabolic with volume growth satisfying (2).
 Theorem 6.11.
@@
 Assume on the contrary that The condition implies that
-in the sense of currents, where and are understood as divisors. Thus, we obtain
+in the sense of currents, where and are understood as divisors. Thus, it yields from that
 In further, it follows from Lemma 6.3 that
@@
 It is therefore
-But, it contradicts with the assumption for This completes the proof. ∎
+which contradicts with the condition (6) for This completes the proof. ∎
 Corollary 6.12.
@@
 Proof.
+It is not hard to show that holds with under the condition for Employing Theorem 6.11, we have holds. Similarly, we have holds with Hence, we prove ∎
 Corollary 6.13 (Five-Value Type Theorem).
@@
 B2 has non-positive sectional curvature
-Assume that has non-positive sectional curvature satisfying (3).
+Assume that has curvature satisfying (3).
 Theorem 6.14.